### v1

# Surface groups among cubulated hyperbolic and one-relator groups

Abstract
Let $X$ be a non-positively curved cube complex with hyperbolic fundamental group. We prove that $\pi_{1}(X)$ has a non-free subgroup of infinite index unless $\pi_{1}(X)$ is either free or a surface group, answering a question of Wise. A similar result for one-relator groups follows, answering a question posed by several authors. The proof relies on a careful analysis of free and cyclic splittings of cubulated groups.
Surface groups – the fundamental groups of closed, aspherical surfaces – and free groups are often seen as close cousins. One way this can be made precise is to look at their subgroup structures. The Nielsen–Schreier theorem asserts that every subgroup of a free group is free, while Whitehead also proved that every subgroup of infinite index in a surface group is free. It is natural to wonder whether this property characterises free and surface groups. Conjectures and questions along these lines have been stated under various hypotheses [4, 3, 20, 18, 55].
To the best of the author’s knowledge, the following very general question is open.
Question 0.1.
Let $G$ be an infinite, finitely presented group, such that every subgroup of infinite index is free. Must $G$ be isomorphic to either a free group or a surface group?
A finitely generated, but infinitely presented, group with every proper subgroup isomorphic to $\mathbb{Z}$ was constructed by Ol’shanskii [36]. Clearly, a group $G$ with a surface subgroup can only have every subgroup of infinite index free if $G$ itself is virtually a surface group, so Question 0.1 can be thought of as a useful precursor to the famous problem of finding surface subgroups.
We provide positive answers to Question 0.1 for two important classes of groups. Recall that a group $G$ is cubulated if $G=\pi_{1}(X)$, where $X$ is a compact, non-positively curved cube complex. Cubulated groups have been a major theme in geometric group theory for several decades, culminating in the work of Agol and Wise on the virtual Haken conjecture [1, 56]. Our main theorem gives a strong positive answer to Question 0.1 for cubulated hyperbolic groups.
Theorem A.
Let $G$ be a cubulated hyperbolic group. Unless $G$ is free or a surface group, $G$ has a one-ended, quasiconvex subgroup of infinite index.
In particular, Theorem A applies to $C^{\prime}(1/6)$ small-cancellation groups [54].
The proof makes essential use of the cubulation hypothesis, and removing it seems to be well beyond current technology. The hyperbolicity hypothesis, on the other hand, can be relaxed to the assumption that $G$ is virtually special in the sense of Haglund–Wise [26]. [^1] Agol’s theorem asserts that cubulated hyperbolic groups are virtually special, so this is indeed a relaxation [1].
Corollary B.
Let $G$ be the fundamental group of a compact, virtually special cube complex. Unless $G$ is free or a surface group, $G$ has a one-ended, cubically convex-cocompact subgroup of infinite index.
Proof.
If $G$ is hyperbolic then the result follows from Theorem A. Otherwise, Caprace–Haglund proved that $G$ contains a cubically convex-cocompact subgroup isomorphic to $\mathbb{Z}^{2}$ [10, Corollary 4], and the result follows unless $\mathbb{Z}^{2}$ has finite index. But, in the latter case, $G$ is a surface group by Bieberbach’s theorem. ∎
This corollary answers a question asked by Wise in the 2014 ICM proceedings [55, Question 13.50].
Mapping tori of free-group automorphisms are another appealing class of groups to which Theorem A applies.
Corollary C.
Let $F$ be a finitely generated free group of rank at least 2. Any semidirect product $G=F\rtimes\mathbb{Z}$ has a finitely generated, one-ended subgroup.
Proof.
Brinkmann proved that $G$ is hyperbolic unless $G$ contains a subgroup isomorphic to $\mathbb{Z}^{2}$ [9]. Since $F$ has rank at least 2, such a subgroup must have infinite index, so the result follows unless $G$ is hyperbolic. In this case, $G$ is cubulated by the work of Hagen–Wise [23, 24]. Since $G$ is neither free nor a surface group, the result again follows, this time by Theorem A. ∎
Historically, Question 0.1 has received the most attention when $G$ is a one-relator group. In this context, the question was attributed to Fine–Rosenberger–Wienke in the Kourovka notebook [18, Question 20.22], but it has also received attention in many other works on one-relator groups, sometimes with additional hypotheses [3, 4, 14, 19]. In this context, it is sometimes referred to as the surface group conjecture. In [52], the author proved the surface group conjecture for Sela’s limit groups, and also for graphs of free groups with cyclic edge groups. [^2] The latter class includes the classes of cyclically pinched and conjugacy pinched one-relator groups considered in [14, 19].
Most recently, the surface group conjecture was addressed by Gardam–Kielak–Logan [20, Conjecture 1.4], who proved it in the case when $G$ has two generators. Combining their work with Theorem A and other recent advances in our understanding of one-relator groups, the surface group conjecture follows.
Theorem D.
Let $G$ be an infinite one-relator group. Unless $G$ is free or a surface group, $G$ has a subgroup of infinite index that is not free.
The alert reader will notice that the conclusion in the case of one-relator groups is slightly weaker than the conclusions for cubulated hyperbolic or free-by-cyclic groups. While Theorem A produces a finitely generated (indeed, quasiconvex) subgroup that is not free, the non-free subgroup produced by Theorem D may be infinitely generated.
Remark 0.2.
In fact, the proof of Theorem D produces a finitely generated non-free subgroup unless $G$ has two generators. In the two-generator case, the stronger result is not true: every finitely generated subgroup of infinite index in the $(1,2)$-Baumslag–Solitar group $BS(1,2)$ is either trivial or isomorphic to $\mathbb{Z}$, but $BS(1,2)$ contains a subgroup isomorphic to the infinitely generated non-free group $\mathbb{Q}_{2}$ of dyadic rationals. (Cf. Question 5.4 below.)
The proof of Theorem A is inspired by the author’s proof of the corresponding statement for graphs of free groups with cyclic edge groups [52]. That proof relied on Whitehead theory to recognise relative free splittings of free groups, so our first task is to generalise that machinery to the context of non-positively curved cube complexes.
Let $X$ be a compact, non-positively curved cube complex. We consider compact, convex subcomplexes $Y$ of the universal cover $\widetilde{X}$, and define a Whitehead complex $\mathrm{Wh}_{X}(Y)$. This is a simplicial complex that encodes the intersections of the bounding hyperplanes of $Y$, and provides a topological model for the complement $\widetilde{X}\smallsetminus Y$. We will be especially interested in the minimal cardinality $k$ of a cut set of vertices of $\mathrm{Wh}_{X}(Y)$: if there is such a cut set, then $Y$ is called a $k$-cut. An essential cube complex $X$ has a 0-cut if and only if $\pi_{1}(X)$ has more than one end.
We analyse the behaviour of the Whitehead complex when $Y$ is cut along a hyperplane. It turns out that $\mathrm{Wh}_{X}(Y)$ decomposes as a connect sum in the natural sense for simplicial complexes, in a way that generalises the work of Manning [34] and Cashen–Macura [13] in the classical context of Whitehead graphs. Our key technical results analyse how cut sets of vertices behave under the connect sum operation. This enables us to relate cuts to links of vertices. The following criterion, a cubical analogue of Whitehead’s cut-vertex lemma, is a sample application.
Theorem E.
Let $X$ be a compact, non-positively curved cube complex. If the link of every vertex is connected and has no cut simplices then $\pi_{1}(X)$ has at most one end.
While 0-cuts detect free splittings, 1-cuts turn out to obstruct analysing splittings over non-trivial subgroups. Fortunately, using recent work of Shepherd [42], as long as $\pi_{1}(X)$ is one-ended we may assume that $X$ has no 1-cuts.
Assuming $\pi_{1}(X)$ is one-ended, our next task is to analyse splittings over $\mathbb{Z}$, and to do this we restrict to the case when $\pi_{1}(X)$ is hyperbolic. The approach generalises work of Cashen–Macura [13] and Cashen [12], who showed that a line pattern in a free group does not split over $\mathbb{Z}$ if and only if there is a bound on the diameters of cut pairs in the Whitehead graph. We prove an analogous result, Theorem 3.11, which asserts that $\pi_{1}(X)$ is cyclically indecomposable if and only if there is a uniform bound on the widths of 2-cuts. The point is that cyclic splittings correspond to periodic 2-cuts, from which a pigeonhole argument implies a bound on the width of a 2-cut if $\pi_{1}(X)$ is cyclically indecomposable. Using the projection-complex machinery of Bestvina–Bromberg–Fujiwara [5], we strengthen this to a uniform bound on the widths of $k$-cuts, independent of $k$.
With these technical results in hand, we are ready to prove Theorem A. Using the Grushko and JSJ decompositions, and since the case of graphs of free groups with cyclic edge groups is already known, we may reduce to the cyclically indecomposable case. By the above analysis, we may assume that $X$ has no 0- or 1-cuts, and that there is a uniform bound $C$ on the widths of $k$-cuts. Agol’s theorem [1] implies that hyperplane stabilisers are separable, so we may pass to a finite-sheeted cover $X_{0}\to X$ with the property that, for some hyperplane $H$ of $\widetilde{X}$, every $\pi_{1}(X_{0})$-translate of $H$ is at distance $>C$ from $H$. Cutting $X_{0}$ along $H$ and passing to a connected component $X^{\prime}$ gives a quasiconvex subgroup that is necessarily of infinite index. But now a 00-cut in $X^{\prime}$ would lead to a $k$-cut in $X$ of width greater than $C$, which is a contradiction, so $\pi_{1}(X^{\prime})$ is not free, as required.
The paper is structured as follows. Section 1 contains preliminaries on non-positively curved complexes, cube complexes and simplicial complexes. Experts may wish to jump straight to Section 2, where we define Whitehead complexes and use them to analyse free splittings, including proving Theorem E. In Section 3 we extend the analysis to cyclic splittings in the hyperbolic case. Section 4 contains the proofs of the main theorems, while some open questions are listed in Section 5.
Acknowledgements
I am grateful to Mark Hagen for useful conversations about this work, most especially for suggesting Lemma 3.16, which greatly simplifies the argument. Thanks also to Macarena Arenas, Jack Button, Chris Cashen, Mark Hagen again, Marco Linton, Lars Louder and Sam Shepherd for comments on an early draft.
1 Preliminaries
Two kinds of cell complexes will be of particular importance in this paper. On the one hand, our main objects of study are non-positively curved cube complexes and their universal covers, which are CAT(0). On the other hand, we will study them by constructing certain simplicial complexes, called Whitehead complexes. In this section, we recall some basic facts about all these classes of complexes. The reader is referred to the standard reference [8] for background material on cube complexes and other metric polyhedral complexes, and for details of much of what follows.
1.1 Non-positively curved cube complxes
A cube complex $X$ is a cell complex constructed from a disjoint union of finite-dimensional unit cubes by identifying faces homeomorphically [8, Example I.7.40(4)].
The link $\mathrm{Lk}_{X}(x)$ of a point $x$ in a metric polyhedral complex $X$ is the space of directions at $x$, i.e. the space of germs of geodesics emanating from $x$ [8, I.7.38], and admits a natural angular metric. When $X$ is a cube complex and $x$ is a vertex, $\mathrm{Lk}_{X}(x)$ has the natural structure of a simplicial complex, where the simplices are the corners of the cubes incident at $x$.
As long as a cube complex $X$ is finite-dimensional, $X$ admits a natural complete geodesic metric in which the interior of every $k$-cube is locally isometric to $\mathbb{R}^{k}$. We will refer to this metric as the $\ell^{2}$-metric on $X$. Gromov’s link condition asserts that a polyhedral complex is non-positively curved – i.e. the metric is locally CAT(0) – if and only if the angular metric on the link of any point is CAT(1) [22, 4.2.A]. Gromov further proved that, when $X$ is a cube complex, the link condition is satisfied if and only if the link of every vertex is flag, meaning that, for every $k\geq 2$, every $k$-clique in the 1-skeleton of $X$ is the boundary of a $k$-simplex [22, 4.2.C].
The features of a non-positively curved cube complex $X$ can often be captured by studying their hyperplanes. Let $C$ be a $k$-dimensional cube, identified with $[-1/2,1/2]^{k}\subseteq\mathbb{R}^{k}$. A hypercube of $C$ is an intersection of $C$ with any coordinate hyperplane of $\mathbb{R}^{k}$. The set of hypercubes of cubes in $X$ form a natural directed system under inclusion, and gluing along these inclusions gives a cube complex with a natural map to $X$. The path components of this cube complexes are the hyperplanes of $X$.
It is immediate from the definition that the natural map from a hyperplane $H$ to $X$ is a local isometry. Therefore, if $X$ is non-positively curved then $H$ is too.
Since hyperplanes are constructed by gluing together hypercubes of $X$, each hyperplane $H\to X$ admits a natural (closed) cubical neighbourhood $N(H)$, which is a bundle over $H$ with fibre a closed interval. Likewise, $N(H)$ has a natural interior $\mathring{N}(H)$, also a bundle over $H$ with fibre an open interval. The hyperplane $H$ is said to be two-sided when these bundles are trivial.
If the map $H\to X$ is injective we say that $H$ is embedded. In this case, the interior $\mathring{N}(H)$ is also naturally embedded in $X$. The complement $X_{0}=X\smallsetminus\mathring{N}(H)$ is a locally convex subcomplex of $X$, and is said to be obtained by cutting $X$ along $H$. In this case, we write $X=X_{0}\cup_{H}$. When $H$ is two-sided, this expresses $X$ as a graph of spaces in the sense of Scott and Wall [39]. As long as $X$ itself is connected, the vertex space $X_{0}$ has at most two components, and if there are are two then we write $X=X_{1}\cup_{H}X_{2}$ in the obvious way.
Finally, we can make use of the following important lemma [8, Proposition II.4.14].
Lemma 1.1.
Let $X,Y$ be complete, connected, geodesic, metric spaces. If $X$ is non-positively curved then any local isometry $Y\to X$ is injective on fundamental groups.
Thus, the inclusion of a hyperplane is $\pi_{1}$-injective, so cutting along $H$ also decomposes the fundamental group $G=\pi_{1}(X)$ as a graph of groups in the sense of Serre [40], with a single edge group $\pi_{1}(H)$.
1.2 CAT(0) cube complexes
When $X$ is non-positively curved, the natural geodesic metric on the universal cover $\widetilde{X}$ is CAT(0), by the Cartan–Hadamard theorem [22, §4]. This gives us access to the standard tools of CAT(0) geometry detailed in [8], which are especially useful when studying a closed, convex subspace $Y\subseteq\widetilde{X}$. For instance, we may define the orthogonal projection $\pi_{Y}:\widetilde{X}\to Y$ as in [8, Proposition II.2.4].
One important consequence of the uniqueness of geodesics in CAT(0) spaces is that any local isometry between CAT(0) spaces is a global isometric embedding [8, Proposition II.4.14]. Thus, the map from a hyperplane $H\to X$ lifts to a convex embedding of the universal cover $\widetilde{H}\hookrightarrow\widetilde{X}$, and so the hyperplanes of $\widetilde{X}$ are the universal covers of the hyperplanes of $X$.
The natural map $N(H)\to X$ is also a local isometry. Thus, the map $N(H)\to H$ also lifts to an embedding
$\widetilde{N(H)}=N(\widetilde{H})\hookrightarrow\widetilde{X}\,.$
The interval bundle $N(\widetilde{H})$ is necessarily trivial because $\widetilde{H}$ is simply connected, so $\widetilde{H}$ is embedded and two-sided, and cutting along $\widetilde{H}$ realises $\widetilde{X}$ as a graph of spaces. Because $\widetilde{X}$ is simply connected it follows that $\widetilde{H}$ separates it into two, and so we have
$\widetilde{X}=\widetilde{U}_{+}\cup_{\widetilde{H}}\widetilde{U}_{-}$
for two convex subcomplexes $\widetilde{U}_{\pm}$. These subcomplexes are called the half-spaces of $\widetilde{H}$.
It is usually important when studying CAT(0) cube complexes to rule out ‘uninteresting’ half-spaces that are within a bounded distance of the corresponding hyperplane.
Definition 1.2.
A hyperplane $\widetilde{H}$ of a CAT(0) cube complex $\widetilde{X}$ is called inessential if some half-space of $\widetilde{H}$ is within a bounded neighbourhood of $\widetilde{H}$. If $\widetilde{X}$ has no inessential hyperplanes then $\widetilde{X}$ itself is called essential. We will call a non-positively curved complex $X$ essential or inessential according to its universal cover.
Fortunately, we may always assume that $X$ is essential by invoking a result of Caprace–Sageev [11, Proposition 3.5].
Lemma 1.3.
Every compact, non-positively curved cube complex $X$ deformation retracts to an essential, convex subcomplex of its first subdivision.
Hyperplanes also make it possible to define a notion of convex hull for a subset of $\widetilde{X}$.
Definition 1.4.
For any subset $Y$ of a CAT(0) cube complex $\widetilde{X}$, the cubical convex hull $\operatorname{Hull}(Y)$ is defined to be the intersection of the half-spaces of $\widetilde{X}$ that contain $Y$.
It is important to be able to control the geometry of convex hulls, and we can do this using a fundamental theorem of Haglund. This theorem makes use of the $\ell^{1}$-metric on (the vertices of) $\widetilde{X}$.
Definition 1.5.
Let $\widetilde{X}$ be a CAT(0) cube complex. The $\ell^{1}$-metric on the 1-skeleton $X^{(1)}$ declares the distance between two points $x$ and $y$ to be equal to the length of the shortest path in $X^{(1)}$ between them.
Note that the $\ell^{1}$ metric on $X^{(1)}$ is also geodesic, and that the $\ell^{1}$ distance between two points is exactly the number of hyperplanes that separate them. The next lemma is useful for relating the $\ell^{1}$ and $\ell^{2}$ distances.
Lemma 1.6.
Let $Y_{1},Y_{2}$ be convex subcomplexes of a CAT(0) cube complex $\widetilde{X}$. Suppose that a geodesic $\gamma$ minimises the distance from $Y_{1}$ to $Y_{2}$. A hyperplane $H$ separates $Y_{1}$ from $Y_{2}$ if and only if $\gamma$ crosses $H$.
Proof.
The necessity is obvious, so it remains to prove that, if $\gamma$ crosses $H$, then $H$ separates $Y_{1}$ from $Y_{2}$. Suppose that $\gamma(0)\in Y_{1}$ and $\gamma$ crosses $H$ at $\gamma(t)$. If $H$ intersects $Y_{1}$ then $\gamma(0)$, $\gamma(t)$ and $\pi_{H}(\gamma(0))$ forms a triangle with two right angles, contradicting the hypothesis that $\widetilde{X}$ is CAT(0). Therefore, $H$ is disjoint from $Y_{1}$, and likewise from $Y_{2}$. Since $\gamma$ crosses $H$ once, it follows that $H$ separates $Y_{1}$ from $Y_{2}$, as required. ∎
We can now state Haglund’s theorem [25, Theorem H]
Theorem 1.7 (Haglund).
Let $\widetilde{X}$ be a $d$-dimensional CAT(0) cube complex and $Y\subseteq\widetilde{X}^{(0)}$ a set of vertices that is $\kappa$-quasiconvex in the $\ell^{1}$-metric. There is $R=R(d,\kappa)$ such that every point of $\operatorname{Hull}(Y)$ is at distance at most $R$ from a point of $Y$.
One useful consequence is to characterise geometrically well-behaved subgroups. A subgroup $H$ of the fundamental group of a non-positively curved cube complex $X$ is called cubically convex-cocompact if $H$ acts cocompactly on a subcomplex $Y\subseteq\widetilde{X}$ that is convex in the $\ell^{2}$ metric.
Corollary 1.8.
Let $X$ be a compact, non-positively curved cube complex. A subgroup $H$ of $\pi_{1}(X)$ is cubically convex-cocompact if and only if the $H$-orbit of some vertex is quasiconvex in the $\ell^{1}$ metric.
1.3 Simplicial complexes
As the above discussion makes clear, our investigations of cube complexes will often be metric. By contrast, when simplicial complexes arise, we will be interested in their combinatorial structure, and especially interested in collections of vertices and simplices that separate. Here, we introduce the necessary terminology.
We already saw that, in a metric complex, the link of a point can be thought of metrically as a space of directions. In a simplicial complex, it is convenient to give an equivalent combinatorial definition, which also extends to certain subcomplexes. A subcomplex $B$ of a simplicial complex $A$ is called full if, whenever all the vertices of a simplex $\sigma$ are in $B$, then $\sigma$ itself is in $B$.
Definition 1.9 (Stars and links).
Let $A$ be a simplicial complex and $B\subseteq A$ a full subcomplex. The open star of $B$, $\mathrm{St}_{A}(B)$, is the union of the set of simplices of $A$ that contain a simplex of $B$ as a face. The link of $B$, $\mathrm{Lk}_{A}(B)$, is the set of faces of simplices of $\mathrm{St}_{A}(B)$ that are disjoint from $B$.
Note that $\mathrm{St}_{A}(B)$ may not be a subcomplex of $A$, but $\mathrm{Lk}_{A}(B)$ always is. Furthermore, after putting a suitable metric on $A$, for any vertex $v$, $\mathrm{Lk}_{A}(v)$ coincides with the metric definition of link given in §1.1.
As mentioned above, we are especially interested in the separation properties of subcomplexes.
Definition 1.10 (Cut sets and connectivity).
Let $A$ be a simplicial complex and $B$ a full subcomplex. Write $A_{B}:=A\smallsetminus\mathrm{St}(B)$. The subcomplex $B$ is said to separate $A$ if $A_{B}$ is disconnected. If $B$ is a simplex we say $B$ is a cut simplex.
In the specific case where $B$ is a set of $k$ pairwise non-adjacent vertices, we say that $B$ is a cut set for $A$. If #$\#B=1$ we say $B$ is a cut vertex, and if #$\#B=2$ we say $B$ is a cut pair. If $A$ has no cut set of cardinality at most $k$, then $A$ is said to be $k$-connected.
2 Free splittings
In 1936, J.H.C. Whitehead explained how to determine whether an element $w$ of a free group $F$ is primitive. For a fixed basis $B$ of $F$, his algorithm uses a graph $\mathrm{Wh}_{B}(w)$ which is now known as the Whitehead graph [51]. His algorithm can equally be interpreted as searching for free splittings $F\cong\langle w\rangle*F^{\prime}$. Many elaborations and generalisations of Whitehead’s argument have been given over the years by, among others, Lyon [33], Stallings [46] and Stong [47]. Diao–Feighn, building on work of Gersten [21], gave an algorithm to compute the Grushko decomposition of a graph of free groups [15]. [^3] See the recent preprint of Dicks for a comprehensive history of this topic [16, §2].
The power of this point of view is that global information about a group $G$ – whether or not $G$ is free, or splits freely, or most generally the structure of the Grushko decomposition of $G$ – is encoded in the local information of a Whitehead graph, which can be read off a presentation. The following theorem, which is the result of combining Whitehead’s algorithm with a theorem of Shenitzer [41, Theorem 5], exemplifies this idea in the case of one-relator groups.
Theorem 2.1.
[…]
Note that $\mathrm{Wh}_{B}(w)$ is exactly the link of the unique vertex in the natural presentation complex for $G$. Our goal is to develop a similar theory for non-positively curved cube complexes. To this end, our first task is to generalise the notion of Whitehead graph to the context of non-positively curved cube complexes.
2.1 Bounding hyperplanes and Whitehead complexes
Let $X$ be a non-positively curved cube complex, let $\widetilde{X}\to X$ be its universal cover and let $Y\subseteq\widetilde{X}$ be a convex subcomplex. Equivalently, we may think of $Y$ as a CAT(0) complex equipped with a locally isometric combinatorial map $Y\to X$ – any such immersion lifts to an embedding $Y\hookrightarrow\widetilde{X}$, and all such embeddings differ by a deck transformation of $X$.
Definition 2.2.
A hyperplane $H$ of $\widetilde{X}$ bounds $Y$ if it is disjoint from $Y$ but no other hyperplane separates $Y$ from $H$. Equivalently, $H$ and $Y$ are disjoint, but $Y$ intersects the (closed) regular neighbourhood $N(H)\cong H\times[0,1]$. We may orient the interval factor so that the intersection is of the form $H_{Y}\times\{0\}$, where $H_{Y}$ is a convex subcomplex of $H$ called the bounding hyperplane component of $Y$ in $H$.
The Whitehead complex associated to such a $Y$ encodes the way that the bounding hyperplanes of $Y$ intersect each other.
Definition 2.3.
Let $X$ be a non-positively curved cube complex, and $Y$ a convex subcomplex of the universal cover $\widetilde{X}$. The Whitehead complex of $Y$, $\mathrm{Wh}_{X}(Y)$, is a simplicial complex defined as follows. The vertices of $\mathrm{Wh}_{X}(Y)$ are the bounding hyperplanes of $Y$. A finite set of bounding hyperplanes =$\mathcal{H}=\{H_{0},\ldots,H_{k}\}$ spans a $k$-simplex $\sigma_{\mathcal{H}}$ in $\mathrm{Wh}_{X}(Y)$ if and only if the intersection $\bigcap_{i=0}^{k}H_{i}$ is non-empty.
Remark 2.4.
The Helly property for hyperplanes of CAT(0) cube complexes, proved by Sageev, asserts that if a set of hyperplanes intersect pairwise then their intersection is non-empty [38, Theorem 4.14]. Hence, Whitehead complexes are flag.
The name is justified by the following example.
Example 2.5.
Let $F$ be a free group, thought of as the fundamental group of a bouquet of circles $R$. (This is equivalent to choosing a basis $B$ for $F$.) The conjugacy class of a non-trivial element $w\in F$ can be represented by an immersion $w:S^{1}\to R$. The mapping cylinder $X$ of this immersion is naturally a non-positively curved square complex. Let * be the one-point space and consider the map $*\to X$ with image equal to the unique vertex of $R$. Then $\mathrm{Wh}_{X}(*)$ is the Whitehead graph $\mathrm{Wh}_{B}(w)$, subdivided once.
More generally, in [13], Cashen and Macura defined generalised Whitehead graphs $\mathrm{Wh}_{B}(T)\{w\}$ associated to a word $w$ in a free group $F$, for any subtree $T$ of the Cayley tree $\widetilde{R}$ of $F$. Our complex $\mathrm{Wh}_{X}(T)$ is the graph obtained from $\mathrm{Wh}_{B}(T)\{w\}$ by subdividing once.
The case when $Y$ is a vertex provides another important example.
Example 2.6.
Consider $\tilde{y}$ a vertex of $\widetilde{X}$ with image $y\in X$. The bounding hyperplanes of $\tilde{y}$ are dual to the edges incident at $y$, and a set of such hyperplanes cross exactly when the edges bound a cube with corner $\tilde{y}$. Therefore, $\mathrm{Wh}_{X}(\tilde{y})\cong\mathrm{Lk}_{X}(y)$ as simplicial complexes.
Whitehead complexes are useful because they provide a topological model for the complement of $Y$: we shall see that $Wh_{X}(Y)$ is homotopy equivalent to the complement $\widetilde{X}\smallsetminus Y$. In order to describe this homotopy equivalence, it is useful to define the link of $Y$, and to extend our discussion of bounding hyperplanes to codimension-$k$ hyperplanes.
Definition 2.7.
Consider a set of $k$ distinct hyperplanes $H_{1},\ldots,H_{k}$ of $Y$ that pairwise intersect. The intersection
$K=\bigcap_{i=1}^{k}H_{i}$
is called a codimension-$k$ hyperplane. It has a closed regular neighbourhood $N(K)\cong K\times[0,1]^{k}$. If the $H_{i}$ are all bounding hyperplanes of $Y$ then $K$ is called a bounding codimension-$k$ hyperplane of $Y$. In this case, we may orient the interval factors so that
$N(K)\cap Y=K_{Y}\times\{0\}^{k}$
for some convex subcomplex $K_{Y}$, called the bounding codimension-$k$ hyperplane component of $K$ in $Y$.
The link of $K_{Y}\times\{0\}^{k}$ in $K_{Y}\times[0,1]^{k}$ consists of unit vectors orthogonal to $K_{Y}\times\{0\}^{k}$, and is naturally of the form $K_{Y}\times\sigma_{K}$, where $\sigma_{K}$ is a $(k-1)$-simplex.
Whenever a codimension-$k$ hyperplane $K$ is contained in a codimension-$l$ hyperplane $L$ (necessarily with $l\geq k$), there are natural inclusions $K_{Y}\subseteq L_{Y}$ and $\sigma_{L}\subseteq\sigma_{K}$, which in turn lead to inclusions
$K_{Y}\times\sigma_{L}\subseteq K_{Y}\times\sigma_{K}\,,\,K_{Y}\times\sigma_{L}\subseteq L_{Y}\times\sigma_{L}\,.$ (1)
Gluing along these inclusions defines the link of the subcomplex $Y$.
Definition 2.8.
The link of a convex subcomplex $Y$ of a CAT(0) cube complex $\widetilde{X}$ is the cell complex $\mathrm{Lk}_{\widetilde{X}}(Y)$ defined from the disjoint union
$\coprod_{K}K_{Y}\times\sigma_{K}$
over all bounding hyperplane components of any codimension, by gluing along the natural inclusions (1). As usual, when $\widetilde{X}$ is the universal cover of a non-positively curved cube complex $X$, we will write $\mathrm{Lk}_{X}(Y)$.
Remark 2.9.
The link $\mathrm{Lk}_{\widetilde{X}}(Y)$ can also be described metrically as the germs of geodesics in $\widetilde{X}$ that are orthogonal to $Y$. Thus, orthogonal projection defines a natural map $\hat{\pi}_{Y}:\widetilde{X}\smallsetminus Y\to\mathrm{Lk}_{\widetilde{X}}(Y)$.
We can now prove that $\mathrm{Wh}_{X}(Y)$ is homotopy equivalent to the complement of $Y$ in two stages.
Lemma 2.10.
For any convex subcomplex $Y$ of a CAT(0) cube complex $\widetilde{X}$, the orthogonal projection $\hat{\pi}_{Y}:\widetilde{X}\smallsetminus Y\to\mathrm{Lk}_{\widetilde{X}}(Y)$ is a homotopy equivalence.
Proof.
As in the remark, we identify the link $\mathrm{Lk}_{\widetilde{X}}(Y)$ with the set of germs of geodesics orthogonal to $Y$. Since $Y$ is a subcomplex, every such germ extends uniquely to a geodesic of length $1$. Therefore, $\hat{\pi}_{Y}$ has a well-defined right-inverse $\rho:\mathrm{Lk}_{\widetilde{X}}(Y)\to\widetilde{X}\smallsetminus Y$ sending the germ of $\gamma$ to $\gamma(1)$. For any $x\notin Y$, let $\gamma_{x}$ denote the unique geodesic from $\hat{\pi}_{Y}(x)$ to $x$. Then
$(x,t)\mapsto\gamma_{x}((1-t)d(x,Y)+t)$
defines a homotopy equivalence between the identity and $\rho\circ\hat{\pi}_{Y}$, which completes the proof. ∎
Finally, we relate links to Whitehead complexes by contracting bounding hyperplane components.
Lemma 2.11.
For any convex subcomplex $Y$ of a CAT(0) cube complex $\widetilde{X}$, there is a homotopy equivalence $\mathrm{Lk}_{\widetilde{X}}(Y)\simeq\mathrm{Wh}_{\widetilde{X}}(Y)$.
Proof.
By definition, $\mathrm{Wh}_{\widetilde{X}}(Y)$ is the complex obtained by contracting each $K_{Y}$ factor to a point in the definition of $\mathrm{Lk}_{\widetilde{X}}(Y)$. Since $\mathrm{Lk}_{\widetilde{X}}(Y)$ is a cell complex and each $K_{Y}$ is contractible, the result follows by, for example, [27, Proposition 0.16]. ∎
The previous two lemmas combine to give the following fundamental result.
Lemma 2.12.
For any convex subcomplex $Y$ of a CAT(0) cube complex $\widetilde{X}$, there is a homotopy equivalence $\widetilde{X}\smallsetminus Y\simeq\mathrm{Wh}_{\widetilde{X}}(Y)$.
Finally, it will be useful to understand the links of vertices in Whitehead complexes. (Note that here, in the setting of simplicial complexes, we mean link in the sense of Definition 1.9.)
Lemma 2.13.
Let $Y$ be a convex subcomplex of a CAT(0) cube complex $\widetilde{X}$ and let $H$ be a bounding hyperplane of $Y$. Then
$\mathrm{Lk}_{\mathrm{Wh}_{\widetilde{X}}(Y)}(H)\cong\mathrm{Wh}_{H}(H_{Y})\,.$
Proof.
Since both simplicial complexes are flag, it suffices to define an isomorphism at the level of 1-skeleta.
The vertices of $\mathrm{Lk}_{\mathrm{Wh}_{\widetilde{X}}(Y)}(H)$ are, by definition, the bounding hyperplanes of $Y$ that cross $H$. For every such hyperplane $K$, the intersection $H\cap K$ is a bounding hyperplane of $H_{Y}$ in $H$. Conversely, every bounding hyperplane of $H_{Y}$ in $H$ extends uniquely to a bounding hyperplane of $Y$ that crosses $H$. Thus, $K\mapsto H\cap K$ defines the required bijection on the 0-skeleta. The bijection extends to an isomorphism of 1-skeleta, because bounding codimension-three hyperplanes of $Y$ contained in $H$ (which correspond to edges of $\mathrm{Lk}_{\mathrm{Wh}_{\widetilde{X}}(Y)}(H)$) are the same thing as bounding codimension-2 hyperplanes of $H_{Y}$ in $H$ (which correspond to edges of $\mathrm{Wh}_{H}(H_{Y})$). ∎
2.2 Ends and splittings
Invoking a famous theorem of Stallings, the number of (Freudenthal) ends of the universal cover of a non-positively curved cube complex $X$ detects the free splittings of the fundamental group [45]. (Note that $\pi_{1}(X)$ is necessarily torsion-free.)
Theorem 2.14 (Stallings).
Let $X$ be a compact, non-positively curved cube complex.
  1. (i)
    $\mathrm{Ends}(\widetilde{X})=0$ if and only if $\pi_{1}(X)$ is trivial.
  2. (ii)
    $\mathrm{Ends}(\widetilde{X})>1$ if and only if $\pi_{1}(X)$ splits freely. [^4] A non-trivial free splitting corresponds to an action on a simplicial tree without a global fixed point and with trivial edge stabilisers. Note that $\mathbb{Z}\cong 1*_{1}$ splits freely.
We will see that the number of ends can be conveniently encoded in the connectivity properties of Whitehead graphs. The next result is prototypical, and deals with the case when $X$ is simply connected.
Proposition 2.15.
Let $X$ be a compact, non-positively curved cube complex. There is a compact CAT(0) cube complex $Y$ mapping to $X$ with $\mathrm{Wh}_{X}(Y)=\varnothing$ if and only if $\pi_{1}(X)$ is trivial.
Proof.
The Whitehead graph $\mathrm{Wh}_{X}(Y)$ is empty if and only if $Y=\widetilde{X}$. Thus such a compact $Y$ exists if and only if $\widetilde{X}$ is compact, meaning that $\pi_{1}(X)$ is finite, i.e. trivial. ∎
In §1.2, we saw that is is often convenient to assume that our cube complexes are essential. In fact, when studying free splittings, we will only need a weaker notion.
Definition 2.16.
A CAT(0) cube complex $\widetilde{X}$ is called freely inessential if some (necessarily compact) hyperplane $H\subseteq\widetilde{X}$ has the property that one of the associated half-spaces is compact. Otherwise, $\widetilde{X}$ is called freely essential.
Remark 2.17.
Any compact CAT(0) cube complex is freely inessential, unless it is a point, and any essential cube complex is freely essential.
The next result explains how Whitehead complexes encode free splittings of the fundamental group.
Proposition 2.18.
Let $\widetilde{X}$ be a freely essential CAT(0) cube complex. There is a compact, convex $Y\subseteq\widetilde{X}$ such that $\mathrm{Wh}_{\widetilde{X}}(Y)$ is non-empty and disconnected if and only if $\widetilde{X}$ has more than one end.
Proof.
Suppose such a $Y$ exists. Since $\mathrm{Wh}_{\widetilde{X}}(Y)$ is disconnected, Lemma 2.12 implies that $\widetilde{X}\smallsetminus Y$ is also disconnected, so $Y$ separates $\widetilde{X}$ into at least two components. If one of these components is bounded then any bounding hyperplane in the corresponding component of $\mathrm{Wh}_{\widetilde{X}}(Y)$ has a compact half space, so $\widetilde{X}$ is freely inessential. Otherwise, all components are unbounded so $\widetilde{X}$ has more than one end.
For the converse direction, suppose that $\widetilde{X}$ has more than one end, so there is a compact subcomplex $Y\subseteq\widetilde{X}$ such that $\widetilde{X}\smallsetminus Y$ has at least two unbounded components. By Theorem 1.7, we may replace $Y$ by its convex hull and assume that $Y$ is convex. Since $\widetilde{X}\smallsetminus Y$ has at least two components, so does $\mathrm{Wh}_{\widetilde{X}}(Y)$ by Lemma 2.12. ∎
The proposition connects Whitehead complexes and free splittings, but the resulting criterion is ineffective because there is no bound on the possible size of a $Y$ for which $\mathrm{Wh}_{X}(Y)$ is disconnected. To make the criterion effective, we need to understand how to build up Whitehead complexes inductively by gluing, generalising the splicing construction of Manning [34] used by Cashen–Macura [13].
2.3 Connected sums of simplicial complexes
The definition of the connected sum of two simplicial complexes is directly analogous to the definition for manifolds.
Definition 2.19 (Connected sum of complexes).
Let $A,B$ be simplicial complexes, let $a\in A$, $b\in B$ be vertices, and let $\phi:\mathrm{Lk}_{A}(a)\to\mathrm{Lk}_{B}(b)$ be an isomorphism. The complex
$A\#_{\phi}B:=A_{a}\cup_{\phi}B_{b}$
is the connected sum of $A$ to $B$ along $\phi$. (Often, we will have a natural identification $v=a=b$, and $\phi$ will be implicit. In this case, we may write $A\#_{v}B$ instead.)
Remark 2.20.
Definition 2.19 generalises the splicing construction used by Manning to analyse Whitehead graphs in finite-index subgroups [34], and exploited further by Cashen and Macura [12, 13]. Up to taking first subdivisions, Manning’s definition coincides with Definition 2.19 when restricted to graphs. (Recall from Example 2.5 that the classical Whitehead graph is the first subdivision of our Whitehead complex.)
The next result explains how cutting $Y$ along a hyperplane decomposes the Whitehead complex as a connected sum. When a convex subcomplex $Y\subseteq\widetilde{X}$ intersects a hyperplane $H$ of $\widetilde{X}$, we will abuse notation by identifying $H$ with $H\cap Y$. Cutting $Y$ along $H$ then leads to the decomposition
$Y=Y_{1}\cup_{H}Y_{2}\,;$
note that $H$ is then a bounding hyperplane of both $Y_{1}$ and $Y_{2}$.
Proposition 2.21 (Cutting and connected sum).
Let $Y$ be a convex subcomplex of a CAT(0) cube complex $\widetilde{X}$. If
$Y=Y_{1}\cup_{H}Y_{2}$
for some hyperplane $H$ of $\widetilde{X}$ then
$\mathrm{Wh}_{X}(Y)\cong\mathrm{Wh}_{X}(Y_{1})\#_{H}\mathrm{Wh}_{X}(Y_{2})\,.$
Proof.
For notational convenience, write $W_{i}=\mathrm{Wh}_{X}(Y_{i})$ for $i=1,2$, and $W=\mathrm{Wh}_{X}(Y)$. Our first task is to define the isomorphism $\mathrm{Lk}_{W_{1}}(H)\cong\mathrm{Lk}_{W_{2}}(H)$ needed to define the connected sum.
A $(k-1)$-simplex $\sigma_{K}$ of $W_{i}$ corresponds to a bounding codimension-$k$ hyperplane $K$ of $Y_{i}$. This simplex $\sigma_{K}$ is contained in $\mathrm{Lk}_{W_{i}}(H)$ exactly when $K$ crosses $H$. A codimension-$k$ hyperplane that crosses $H$ bounds $Y_{1}$ if and only it bounds $Y_{2}$, so $\mathrm{Lk}_{W_{1}}(H)$ is naturally identified with $\mathrm{Lk}_{W_{2}}(H)$ as required. Let $L$ denote the natural copy of $\mathrm{Lk}_{W_{1}}(H)\equiv\mathrm{Lk}_{W_{2}}(H)$ in $W_{1}\#_{H}W_{2}$.
[…]
  1. (i)
    simplices $\sigma_{K}$ in $L$ correspond to bounding codimension-$k$ hyperplanes of $Y$ that cross $H$;
  2. (ii)
    simplices $\sigma_{K}$ of $W_{1}$ that are not adjacent to $H$ correspond to bounding codimension-$k$ hyperplanes of $Y_{1}$ disjoint from $H$;
  3. (iii)
    simplices $\sigma_{K}$ of $W_{2}$ that are not adjacent to $H$ correspond to bounding codimension-$k$ hyperplanes of $Y_{2}$ disjoint from $H$.
Since this list also covers all the codimension-$k$ hyperplanes of $Y$, and inclusions are preserved, the result follows. ∎
2.4 Cut-set lemmas in connnected sums
To understand splittings of fundamental groups of non-positively curved cube complexes, the key idea is to combine Proposition 2.21 with an understanding of how the connectivity properties of simplicial complexes, in the sense of Definition 1.10, change under connected sum. The first lemma is prototypical.
Lemma 2.22.
Suppose that $A$ and $B$ are simplicial complexes. If $B$ is connected but $A\#_{v}B$ is disconnected, then $v$ is a cut vertex in $A$.
Proof.
Let $L$ denote the link of $v$ and $C(L)$ the cone on $L$. Then $B=B_{v}\cup_{L}C(L)$ so, because $C(L)$ is contractible, the Mayer–Vietoris sequence for reduced homology gives that
$\tilde{H}_{0}(L)\to\tilde{H}_{0}(B_{v})\to\tilde{H}_{0}(B)\to 0$
is exact. Therefore, $\tilde{H}_{0}(L)$ surjects $\tilde{H}_{0}(B_{v})$ because $B$ is connected.
On the other hand, a second application of Mayer–Vietoris to $A\#_{v}B=A_{a}\cup_{L}B_{b}$ gives that
$\tilde{H}_{0}(L)\to\tilde{H}_{0}(A_{v})\oplus\tilde{H}_{0}(B_{v})\to\tilde{H}_{0}(A\#_{v}B)\to 0$
is exact, so $\tilde{H}_{0}(L)$ does not surject $\tilde{H}_{0}(A_{v})\oplus\tilde{H}_{0}(B_{v})$, because $A\#_{v}B$ is disconnected. Hence $\tilde{H}_{0}(A_{v})\neq 0$, i.e. $v$ is a cut vertex in $A$. ∎
Since we will need to apply the lemma inductively, we also need to understand when cut vertices can arise, and more generally cut simplices.
Lemma 2.23.
Suppose that $A$ and $B$ are connected and a simplex $\sigma$ of $A$ is a cut simplex of $A\#_{v}B$. Then one of the following holds:
  1. (i)
    $v$ is a cut vertex in $B$;
  2. (ii)
    $\sigma$ is a cut simplex in either $A$ or $B$; or
  3. (iii)
    $\sigma$ is contained in the link of $v$, and the simplex $\tau=\langle\sigma,v\rangle$ spanned by $\sigma$ and $v$ together is a cut simplex in both $A$ and $B$.
In particular, if $A\#_{v}B$ has a cut simplex then either $A$ or $B$ has a cut simplex.
Proof.
Again, let $C=A\#_{v}B$ for notational convenience. The hypothesis that $\sigma$ is a cut simplex of $C$ means that $C_{\sigma}$ is disconnected. The proof now divides into cases, according to whether $\sigma$ is in $A$ but not in $B$, or $\sigma$ is contained in the canonical copy of the link of $v$.
In the first case, when $\sigma$ is in $A$ but not in $B$, we have
$C_{\sigma}=A_{\sigma}\#_{v}B\,,$
so Lemma 2.22 implies that $v$ is a cut vertex in $B$, as in item (i).
In the second case, if $\sigma$ is in the link of $v$ then
$C_{\sigma}=A_{\sigma}\#_{v}B_{\sigma}\,,$
so Lemma 2.22 leads to two further subcases: either one of $A_{\sigma}$ or $B_{\sigma}$ is disconnected, or $v$ is a cut vertex in both $A_{\sigma}$ and $B_{\sigma}$. The former corresponds to item (ii), while the latter case gives item (iii). ∎
The above results will help us to understand free splittings of $\pi_{1}(X)$. More generally, we will need to understand how arbitrary cut sets decompose.
Lemma 2.24.
Suppose that $U$ is a full set of vertices of $A$ and $V$ is a full set of vertices of $B$, both non-adjacent to a common vertex $v$. If $U\cup V$ is a cut set of $A\#_{v}B$ then: either
  1. (i)
    $V$ is a cut set in $B$; or
  2. (ii)
    $U\cup\{v\}$ is a cut set in $A$.
Proof.
Again, let $C=A\#_{v}B$ for notational convenience. Because $v$ is adjacent to neither $U$ nor $V$,
$C_{U\cup V}=A_{U}\#_{v}B_{V}$
which is disconnected by hypothesis. Therefore, by Lemma 2.22, either $B_{V}$ is disconnected, giving item (i), or $v$ is a cut vertex in $A_{U}$, giving item (ii). ∎
2.5 Cuts and the generalised Whitehead’s lemma
In this section, we combine the results developed so far to prove a sufficient condition for the fundamental group of a compact, non-positively curved cube complex $X$ to be one-ended. These conditions are phrased in terms of $k$-cuts.
Definition 2.25 ($k$-cuts).
Consider a finite CAT(0) cube complex equipped with an immersion $Y\to X$. If $\mathrm{Wh}_{X}(Y)$ has a finite cut set then $Y$ is said to be a cut. More precisely, if $\mathrm{Wh}_{X}(Y)$ is $(k-1)$-connected but not $k$-connected – i.e. if the smallest cut set for $\mathrm{Wh}_{X}(Y)$ has cardinality $k$ – then $Y$ is said to be a $k$-cut. If $\{H_{1},\ldots,H_{k}\}$ is a minimal cut set for a $k$-cut $Y$ then we say that $Y$ is a cut between $H_{1},\ldots,H_{k}$.
Our analysis of splittings of fundamental groups of non-positively curved cube complexes will rest on understanding cuts. We start with 0-cuts, which characterise free splittings. The results of §2.4 explain how 0-cuts decompose.
Lemma 2.26.
Consider a finite CAT(0) cube complex equipped with a locally isometric immersion $Y\to X$, and let $Y=Y_{1}\cup_{H}Y_{2}$ for some hyperplane $H$ of $Y$. If $Y$ is a 0-cut then either one of $Y_{1}$ or $Y_{2}$ is a 0-cut, or both $Y_{1}$ and $Y_{2}$ are 1-cuts.
Proof.
The immersion lifts to an embedding of $Y$ as a convex subcomplex of the universal cover $\widetilde{X}$, and the hyperplane $H$ of $Y$ then extends uniquely to a hyperplane of $\widetilde{X}$, still denoted by $H$. Proposition 2.21 implies that
$\mathrm{Wh}_{X}(Y)\cong\mathrm{Wh}_{X}(Y_{1})\#_{H}\mathrm{Wh}_{X}(Y_{2})\,,$
so the result follows from Lemma 2.22. ∎
To iterate the argument, we also need to analyse cut simplices.
Lemma 2.27.
Consider a finite CAT(0) cube complex equipped with a locally isometric immersion $Y\to X$, and let $Y=Y_{1}\cup_{H}Y_{2}$ for some hyperplane $H$ of $Y$. If $\mathrm{Wh}_{X}(Y)$ has a cut simplex, then one of $\mathrm{Wh}_{X}(Y_{1})$ or $\mathrm{Wh}_{X}(Y_{2})$ has a cut simplex.
Proof.
This follows in the same way as Lemma 2.26, using Proposition 2.21 and Lemma 2.23. ∎
We now have all the tools needed to prove Theorem E, which we think of as a version of Whitehead’s lemma for non-positively curved cube complexes.
Proof of Theorem E.
Let $Y$ be a CAT(0) cube complex, locally isometrically immersed in $X$. We prove that $\mathrm{Wh}_{X}(Y)$ is connected without cut simplices, by induction on the number of vertices of $Y$.
In the base case, $Y$ is a single vertex $y$, so $\mathrm{Wh}_{X}(Y)\cong\mathrm{Lk}_{X}(y)$ by Example 2.6 and there is nothing to prove. If $Y$ has more than one vertex, we may choose a (necessarily separating) hyperplane $H$ of $Y$ and write $Y=Y_{1}\cup_{H}Y_{2}$. By the inductive hypothesis, both $Y_{1}$ and $Y_{2}$ are connected without cut simplices, and therefore so is $Y$ by Lemma 2.27.
In conclusion, $X$ has no 0-cuts, so Proposition 2.18 implies that $\pi_{1}(X)$ has at most one end. ∎
Theorem E provides a convenient certificate that $\pi_{1}(X)$ does not split freely. However, the necessary condition that it provides is not sufficient: there are non-positively curved square complexes in which the link of a vertex contains a cut vertex, and yet the universal cover is one-ended. The following example is based on [28, Example 2.1].
Example 2.28.
Consider the rank-two free group $F=\langle a,b\rangle$, thought of as the fundamental group of the wedge of two circles $\Gamma=S^{1}\vee S^{1}$. Consider the element $w=abab^{2}ab^{3}$ of $F$ and realise the corresponding cyclic word by an immersion of a circle $\gamma:S^{1}\to\Gamma$. Let $X$ be the non-positively curved square complex obtained by gluing either end of a cylinder $S^{1}\times[0,1]$ to two copies of $\Gamma$, using $\gamma$ as the attaching map at each end. The links of the two vertices of $X$ are both isomorphic to the Whitehead graph of $w$, which has a cut vertex.
The fundamental group of $X$ is the double
$\pi_{1}(X)=F*_{\langle w\rangle}F$
by the Seifert–van Kampen theorem. Applying the automorphism $\phi$ of $F_{2}$ setting $\phi(a)=ab^{-2}$ and $\phi(b)=b$ sends $w\mapsto w^{\prime}=ab^{-1}a^{2}b$. We may realise $w^{\prime}$ by an immersion $\gamma^{\prime}:S^{1}\to\Gamma$ and glue along a cylinder as before to obtain a new square complex $X^{\prime}$ with $\pi_{1}(X^{\prime})\cong\pi_{1}(X)$. However, the links of the two vertices of $X^{\prime}$ are isomorphic to the Whitehead graph of $w^{\prime}$, which consists of two triangles glued along an edge, and in particular is connected with no cut vertices. Therefore, by Theorem E, $\pi(X)\cong\pi_{1}(X^{\prime})$ is one-ended, even though the links of $X$ have cut vertices.
In the two-dimensional case, a square complex $X$ can always be modified so that the sufficient condition of Theorem E can be applied. A map of square complexes $\phi:X^{\prime}\to X$ is called an unfolding of $X$ if (i) the restriction of $\phi$ to 1-skeleta is a combinatorial homotopy equivalence of graphs; and (ii) $\phi$ is a homeomorphism on the complements of the 1-skeleta.
Proposition 2.29.
Let $X$ be a compact, non-positively curved square complex. There is an unfolding $\phi:X^{\prime}\to X$ such that
$X^{\prime}=\Gamma\vee X_{1}\vee\ldots\vee X_{m}\vee S_{1}\vee\ldots\vee S_{n}$
is also a compact, non-positively curved square complex and, furthermore:
  1. (i)
    $\Gamma$ is a graph;
  2. (ii)
    each $X_{i}$ has the property that every link is connected without cut vertices or cut edges;
  3. (iii)
    each $S_{j}$ is a square.
Proof.
The existence of $X^{\prime}$ is [53, Lemma 4.6] stated for square complexes (see also [31, §3]); it remains to prove that $X^{\prime}$ is non-positively curved. To do this, it suffices to check Gromov’s link condition, that any cycle $\gamma$ in a link of a vertex of $X^{\prime}$ should have length at least $2\pi$ in the angular metric [8, Lemma II.5.6]. Because $\phi$ is a homeomorphism on the complement of the 1-skeleton, the induced maps on links are injective on edges and hence immersions. Thus, $\gamma$ is also an immersed cycle in a link of $X$, so has length at least $2\pi$ by the link condition for $X$. ∎
The square complexes $X_{i}$ provided by the proposition satisfy the hypotheses of Theorem E. Thus, the wedge decomposition $X^{\prime}$ actually gives the Grushko decomposition of $X$:
$\pi_{1}(X)\cong\pi_{1}(\Gamma)*\pi_{1}(X_{1})*\ldots*\pi_{1}(X_{m})\,,$
and each $\pi_{1}(X_{i})$ is non-trivial and freely indecomposable. Furthermore, the unfolding procedure that produces $X^{\prime}$ from $X$ is algorithmic: one simply unfolds an edge of the 1-skeleton whenever the link of a vertex contains a cut vertex. In summary, Theorem E and Proposition 2.29 together provide an effective, geometric method to determine the Grushko decomposition of the fundamental group of a non-positively curved square complex.
2.6 Shepherd’s theorem and $1$-cuts
The results of §2.5 show that 0-cuts detect when an essential cube complex has more than one end, but they also illustrate the fact that the existence of 1-cuts makes it difficult to determine whether or not $\pi_{1}(X)$ splits. Therefore, for one-ended $\pi_{1}(X)$, we would like to be able to choose $X$ to have no 1-cuts. In the 2-dimensional case, this can be done by unfolding the 1-skeleton as in Proposition 2.29. To handle the general case, however, we can use a recent deep theorem of Shepherd.
A quarter-space of a CAT(0) cube complex is the intersection of two half-spaces. A quarter-space is said to be deep if it is not contained in a bounded neighbourhood of the intersection of its two defining hyperplanes. The next result is [42, Theorem 1.5].
Theorem 2.30 (Shepherd’s theorem).
If $G$ is one-ended and the fundamental group of a non-positively curved cube complex, then $G=\pi_{1}(X)$ for an essential non-positively curved cube complex $X$, such that the universal cover $\widetilde{X}$ has the property that all half-spaces of $\widetilde{X}$ are one-ended and all quarter-spaces are deep.
The following lemma connects Shepherd’s theorem to 1-cuts.
Lemma 2.31.
Let $\widetilde{X}$ be a one-ended, essential CAT(0) cube complex in which all quarter-spaces are deep. If $\widetilde{X}$ has a 1-cut, then some half-space of $\widetilde{X}$ has more than one end.
Proof.
Let $Y\subseteq\widetilde{X}$ be a 1-cut, i.e. $Y$ is compact and convex and $\mathrm{Wh}_{\widetilde{X}}(Y)$ has a cut vertex $H$. Let $U$ be the half-space of $H$ that contains $Y$.
We claim that $U$ is freely essential. Indeed, suppose that some hyperplane $K$ of $U$ had a compact half-space $V$. If $K$ were disjoint from $H$ then $K$ would also be an inessential hyperplane in $\widetilde{X}$, contradicting the hypothesis that $\widetilde{X}$ is essential. If $K$ crossed $H$ then $V$ would be a compact quarter-space in $\widetilde{X}$, contradicting the hypothesis that all quarter-spaces are deep.
Now, $\mathrm{Wh}_{U}(Y)$ is obtained from $\mathrm{Wh}_{\widetilde{X}}(Y)$ by deleting the star of $H$, so is disconnected. Therefore, $U$ has more than one end by Proposition 2.18, as required. ∎
Combining Shepherd’s theorem with the lemma, we may always assume that one-ended cube complexes have no 1-cuts.
Corollary 2.32.
If $G$ is the fundamental group of a compact, non-positively curved cube complex and $G$ is one-ended, then $G=\pi_{1}(X)$, where $X$ is a compact, essential, non-positively curved cube complex without 0-cuts or 1-cuts.
3 Cyclic splittings
Having understood the free splittings of $\pi_{1}(X)$, we may henceforth assume that $\pi_{1}(X)$ is one-ended. Our next task is to adapt the analysis of the previous section to understand splittings of $\pi_{1}(X)$ over $\mathbb{Z}$. We restrict attention to the case where $\pi_{1}(X)$ is hyperbolic and, for brevity, we will say hyperbolic cube complex to mean a compact, connected, non-positively curved cube complex $X$ with $\widetilde{X}$ (and hence $\pi_{1}(X)$) Gromov-hyperbolic. Let $\delta$ be a hyperbolicity constant for $\widetilde{X}$.
Cashen–Macura studied line patterns in free groups, i.e. relatively hyperbolic structures on free groups with cyclic parabolics [13]. They explained how to recognise cut pairs in Otal’s decomposition space – i.e. the Bowditch boundary – by analysing cut pairs in generalised Whitehead graphs. Cashen went on to use this analysis to recognise relative cyclic splittings [12]. Barrett used similar ideas to describe an algorithm that computes the Bowditch JSJ decomposition of any hyperbolic group [2].
Our first task is to adapt their ideas to the setting of hyperbolic cube complexes.
3.1 Periodic $2$-cuts
Suppose that $\pi_{1}(X)$ is one-ended. If every splitting of $\pi_{1}(X)$ with cyclic edge groups is trivial, then $\pi_{1}(X)$ is said to be cyclically indecomposable. The main theorem of this section is a combinatorial characterisation of cyclic indecomposibility. Just as we understood free splittings of $\pi(X)$ in terms of 0-cuts and 1-cuts of $X$, so cyclic splittings can be understood using 2-cuts.
As we will see, cyclic splittings give rise to 2-cuts, but the converse is not quite true: in order to construct a cyclic splitting from a 2-cut, we also need to consider the combinatorial types of the bounding hyperplane components.
Definition 3.1 (Abstract hyperplane component).
Let $X$ be compact, non-positively curved cube complex. An abstract hyperplane component over $X$ consists of the following data:
  1. (i)
    a compact, convex subcomplex $Z$ of a hyperplane $H\subseteq\widetilde{X}$;
  2. (ii)
    a non-trivial partition of $\pi_{0}(\mathrm{Wh}_{H}(Z))$.
Two abstract hyperplane components $Z\subseteq H$ and $Z^{\prime}\subseteq H^{\prime}$ are of isomorphic type if there is a deck transformation $\phi\in\pi_{1}(X)$ such that $\phi(Z)=Z^{\prime}$, inducing an isomorphism $\mathrm{Wh}_{H}(Z)\cong\mathrm{Wh}_{H^{\prime}}(Z^{\prime})$ that preserves the chosen partitions of $\pi_{0}$.
An orientation of an abstract hyperplane component consists of the following extra datum:
  1. (iii)
    a choice of one of the two half spaces of the hyperplane $H$.
Two oriented abstract hyperplane components $Z\subseteq H$ and $Z^{\prime}\subseteq H^{\prime}$ are of opposite type if they are of isomorphic type and, furthermore, the isomorphism $\phi$ can be chosen to reverse the chosen orientations.
This definition enables us to define a periodic 2-cut.
Definition 3.2 (Periodic 2-cut).
Let $X$ be a compact, non-positively curved cube complex. Consider a 2-cut $Y$ in $X$ between hyperplanes $H,H^{\prime}$. Define an oriented abstract hyperplane component corresponding to $H$ as follows:
  1. (i)
    $Z=H_{Y}$;
  2. (ii)
    the partition on $\pi_{0}(\mathrm{Wh}_{H}(Z))=\pi_{0}(\mathrm{Lk}_{\mathrm{Wh}_{X}(Y)})(H)$ is induced by connectivity in $\mathrm{Wh}_{X}(Y)_{\{H,H^{\prime}\}}$;
  3. (iii)
    the orientation is determined by the half-space of $H$ that contains $Y$.
Define an oriented abstract hyperplane component corresponding to $H^{\prime}$ in the same way. The 2-cut $Y$ is said to be periodic if the two oriented abstract hyperplane components are of opposite type. In this case, the element $\phi\in\pi_{1}(X)$ sending $H\to H^{\prime}$ is called the associated covering transformation of $Y$.
We shall see that periodic 2-cuts characterise cyclic splittings. However, first we need to relate cyclic splittings to Whitehead complexes.
Lemma 3.3.
Let $X$ be an essential hyperbolic cube complex with $\pi_{1}(X)$ one-ended. Then $\pi_{1}(X)$ splits over $\mathbb{Z}$ if and only if some infinite cyclic subgroup acts cocompactly on some isometrically embedded subcomplex $\widetilde{Z}$ with $\mathrm{Wh}_{X}(\widetilde{Z})$ disconnected.
Proof.
By the algebraic annulus theorem of Dunwoody–Swenson [17], and noting that $\pi_{1}(X)$ is torsion-free, $\pi_{1}(X)$ splits over a $\mathbb{Z}$ subgroup if and only if some subgroup $C\cong\mathbb{Z}$ satisfies that the number of relative ends $e(G,C)>1$. Because $\pi_{1}(X)$ is hyperbolic, $C$ is quasiconvex.
Choosing a base vertex * in $\widetilde{X}$, the condition that $e(G,C)>1$ means that, for some $R$, the $R$-neighbourhood of the orbit $C*$ separates $\widetilde{X}$ into at least two infinite components. Let $\widetilde{Z}$ be the cubical convex hull of the $R$-neighbourhood of $C*$, which still separates $\widetilde{X}$ into at least two infinite components. By Haglund’s theorem (Theorem 1.7), $C$ acts cocompactly on $\widetilde{Z}$. Because $\widetilde{Z}$ separates $\widetilde{X}$ into two components, $\mathrm{Wh}_{X}(\widetilde{Z})$ is disconnected by Lemma 2.12, as required.
Conversely, suppose that some $C\cong\mathbb{Z}$ acts cocompactly on some convex subcomplex $\widetilde{Z}$ with $\mathrm{Wh}_{X}(\widetilde{Z})$ disconnected, so $\widetilde{Z}$ separates $\widetilde{X}$ into two components by Lemma 2.12. Either some complementary component remains within a bounded distance of $\widetilde{Z}$, contradicting the hypothesis that $X$ is essential, or all complementary components are unbounded away from $\widetilde{Z}$, meaning that $e(G,C)>1$, as required. ∎
Lemma 3.3 makes it easy to prove that periodic 2-cuts characterise cyclic splittings.
Theorem 3.4 (Cyclic splittings and periodic 2-cuts).
Let $X$ be an essential hyperbolic cube complex. Suppose that $\pi_{1}(X)$ is one-ended, and suppose furthermore that $X$ has no 0- or 1-cuts. Then $\pi_{1}(X)$ is cyclically decomposable if and only if there is a periodic 2-cut $Y\to X$.
Proof.
By Lemma 3.3, $\pi_{1}(X)$ is cyclically decomposable if and only if there is a convex $\mathbb{Z}$-periodic subcomplex $\widetilde{Z}\subseteq\widetilde{X}$ that separates $\widetilde{X}$. By Lemma 2.12, $\widetilde{Z}$ separates $\widetilde{X}$ if and only if $\mathrm{Wh}_{X}(\widetilde{Z})$ is disconnected.
Suppose that there is a periodic 2-cut $Y\subseteq\widetilde{X}$ between hyperplanes $H,H^{\prime}$. By definition, $H,H^{\prime}$ define a cut pair in $\mathrm{Wh}_{X}(Y)$, and the corresponding oriented abstract hyperplane components have opposite type. Therefore, we may define the amalgamated subcomplex
$Y\cup_{\phi(H)}\phi(Y)\,.$
Iterating using all powers of $\phi$, we may set $\widetilde{Z}$ to be the infinite amalgamation
$\cdots\cup_{\phi^{-1}(H)}\phi^{-1}(Y)\cup_{H}Y\cup_{\phi(H)}\phi(Y)\cup_{\phi^{2}(H)}\cdots\,,$
which is $\langle\phi\rangle$-invariant and cocompact. By Proposition 2.21, $\mathrm{Wh}_{X}(\widetilde{Z})$ is constructed by iteratively splicing copies of $\mathrm{Wh}_{X}(Y)$ along copies of $H$, meaning that $\mathrm{Wh}_{X}(\widetilde{Z})$ is the infinite union
$\cdots\cup_{L}\mathrm{Wh}_{X}(Y)_{\{H,H^{\prime}\}}\cup_{L}\mathrm{Wh}_{X}(Y)_{\{H,H^{\prime}\}}\cup_{L}\mathrm{Wh}_{X}(Y)_{\{H,H^{\prime}\}}\cup_{L}\cdots\,,$
where $L=\mathrm{Lk}_{\mathrm{Wh}_{X}(Y)}(H)\cong\mathrm{Wh}_{H}(H_{Y})$ by Lemma 2.13. The partitions that are part of the data of an abstract hyperplane component imply that there is a class $c\in\tilde{H}_{0}(L)$ with the property that $c$ and $\phi_{*}(c)$ have the same image in $\tilde{H}_{0}(\mathrm{Wh}_{X}(Y)_{\{H,H^{\prime}\}})$. Applying Mayer–Vietoris and using that homology commutes with direct limits, it follows that $\tilde{H}_{0}(\mathrm{Wh}_{X}(\widetilde{Z}))$ is non-trivial, so $\mathrm{Wh}_{X}(\widetilde{Z})$ is indeed disconnected as required.
Conversely, suppose that there is a convex $\mathbb{Z}$-periodic subcomplex $\widetilde{Z}\subseteq\widetilde{X}$ with $\mathrm{Wh}_{X}(\widetilde{Z})$ disconnected. The compact quotient $Z=\mathbb{Z}\backslash\widetilde{Z}$ has fundamental group $\mathbb{Z}$, and therefore $H^{1}(Z)\cong\mathbb{Z}$. Since the first cohomology is generated by cocycles dual to hyperplanes, some hyperplane $H$ of $Z$ is crossed by any loop representing the generator $\phi$ of $\mathbb{Z}$. Since $\pi_{1}(H)$ is contained in the kernel of the dual map
$\mathbb{Z}\cong\pi_{1}(Z)\to\mathbb{Z}$
it follows that $H$ is simply connected, and so lifts to a hyperplane (still denoted by $H$) of $\widetilde{Z}$. By [11, Lemma 2.3], after replacing $\phi$ by a proper power we may assume that $\phi(H)$ is disjoint from $H$.
Let $\widetilde{Z}_{\pm}$ be the two half-spaces of $H$, so
$\widetilde{Z}=\widetilde{Z}_{-}\cup_{H}\widetilde{Z}_{+}\,.$
We claim that $\mathrm{Wh}_{X}(\widetilde{Z}_{+})$ and $\mathrm{Wh}_{X}(\widetilde{Z}_{-})$ are connected. Indeed, $\widetilde{Z}_{+}$ may be written as the ascending union of compact, convex subcomplexes $Y_{n}$ where
$Y_{n}=Y\cup_{\phi(H)}\phi(Y)\cup_{\phi^{2}(H)}\ldots\cup_{\phi^{n-1}(H)}\phi^{n-1}(Y)$
and, correspondingly, the Whitehead complex of $\widetilde{Z}_{+}$ may may be written as an ascending union
$\mathrm{Wh}_{X}(\widetilde{Z}_{+})=\bigcup_{n\geq 1}\mathrm{Wh}_{X}(Y_{n})_{\phi^{n}(H)}\,.$
Thus, if $\mathrm{Wh}_{X}(\widetilde{Z}_{+})$ were disconnected then it would follow that $\mathrm{Wh}_{X}(Y_{n})_{\phi^{n}(H)}$ were disconnected for some $n$. If so then $Y_{n}$ would be a 1-cut in $X$, contradicting the assumption. Thus, $\mathrm{Wh}_{X}(\widetilde{Z}_{+})$ is connected and, symmetrically, so is $\mathrm{Wh}_{X}(\widetilde{Z}_{-})$. Since
$\mathrm{Wh}_{X}(\widetilde{Z})=\mathrm{Wh}_{X}(\widetilde{Z}_{-})\#_{H}\mathrm{Wh}_{X}(\widetilde{Z}_{+})$
is disconnected, it follows that $H$ is a cut vertex in both $\mathrm{Wh}_{X}(\widetilde{Z}_{-})$ and $\mathrm{Wh}_{X}(\widetilde{Z}_{+})$ by Lemma 2.22.
Now, $H^{\prime}=\phi(H)$ is a hyperplane of $\widetilde{Z}_{+}$ disjoint from $H$, and cutting along $H^{\prime}$ decomposes $\widetilde{Z}_{+}$ as
$\widetilde{Z}_{+}=Y\cup_{H^{\prime}}\phi(\widetilde{Z}_{+})$
where $Y=\widetilde{Z}_{+}\cap\phi(\widetilde{Z}_{-})$; note that $H$ is a bounding hyperplane of $Y$. Hence,
$\mathrm{Wh}_{X}(\widetilde{Z}_{+})=\mathrm{Wh}_{X}(Y)\#_{H^{\prime}}\mathrm{Wh}_{X}(\phi(\widetilde{Z}_{+}))$
and so, since we saw above that $\mathrm{Wh}_{X}(\phi(\widetilde{Z}_{+}))\cong\mathrm{Wh}_{X}(\widetilde{Z}_{+})$ is connected, $Y$ is a 2-cut between $H$ and $H^{\prime}$ by Lemma 2.24. Finally, note that $Y$ is a periodic 2-cut by construction. ∎
3.2 A bound on the widths of $2$-cuts
Although Theorem 3.4 provides a useful characterisation of the existence of cyclic splittings, in order to make it effective we need an a priori bound on the complexity of the 2-cuts that arise. Such a bound is the main result of this section, and is proved in Theorem 3.11 below. It is phrased in terms of the width of a 2-cut.
Definition 3.5 (Width of a $k$-cut).
Let $X$ be a compact, non-positively curved cube complex, and consider a $k$-cut $Y\to X$ between bounding hyperplanes $\{H_{1},\ldots,H_{k}\}$ for $k\geq 2$. The width of $Y$ is the shortest $\ell^{2}$-distance between any distinct $H_{i}$ and $H_{j}$.
The next lemma provides some geometric control on cuts, assuming that there are no 1-cuts.
Lemma 3.6.
Let $\widetilde{X}$ be a one-ended a CAT(0) cube complex without 0- or 1-cuts. Consider a $k$-cut $Y$ between bounding hyperplanes $H_{1},\ldots,H_{k}$. Then $Y\cap\mathrm{Hull}\left(\bigcup_{i=1}^{k}H_{i}\right)$ is also a $k$-cut between $H_{1},\ldots,H_{k}$.
Proof.
Let $K_{1},\ldots K_{l}$ be the hyperplanes of $Y$ that have a half-space $U_{j}$ such that $H_{i}\subseteq U_{j}$ for all $i$. Write $Y_{n}=Y\cap\bigcap_{j=1}^{n}U_{j}$ for each $0\leq n\leq l$. Note that
$Y_{l}=Y\cap\mathrm{Hull}\left(\bigcup_{i=1}^{k}H_{i}\right)\,.$
To prove the lemma, we show by induction on $n$ that each $Y_{n}$ is a $k$-cut.
In the base case, $Y_{0}=Y$ and there is nothing to prove. For the inductive step, let $n\geq 1$. If $K_{n}$ is disjoint from $Y_{n-1}$ then $Y_{n-1}=Y_{n}$ and there is, again, nothing to prove. Otherwise, $K_{n}$ decomposes $Y_{n-1}$ as
$Y_{n-1}=Y_{n}\cup_{K_{n}}Y^{\prime}_{n}$
where $Y_{n}=Y_{n-1}\cap U_{n}$. Then $H_{1},\ldots,H_{k}$ all bound $Y_{n}$ and not $Y^{\prime}_{n}$ (since they are disjoint from $K_{n}$, and $K_{n}$ cannot separate $Y$ from any $H_{i}$). By Proposition 2.21
$\mathrm{Wh}_{X}(Y_{n-1})=\mathrm{Wh}_{X}(Y_{n})\#_{K_{n}}\mathrm{Wh}_{X}(Y^{\prime}_{n})\,,$
and $K_{n}$ is not a cut vertex in $\mathrm{Wh}_{X}(Y^{\prime}_{n})$ because $\widetilde{X}$ has no 1-cuts. Therefore $\{H_{1},\ldots,H_{k}\}$ is a cut set in $\mathrm{Wh}_{X}(Y_{n})$ by Lemma 2.24, so $Y_{n}$ is also a $k$-cut between $H_{1},\ldots,H_{k}$. Taking $n=l$, the result follows. ∎
We now turn our attention to some CAT(0) geometry of hyperplanes. Consider a 2-cut $Y$ between hyperplanes $H_{1}$ and $H_{2}$, and let $\gamma$ be an $\ell^{2}$ geodesic in between $H_{1}$ and $H_{2}$. In order to prove the desired bound, we need to find long chains of mutually disjoint hyperplanes that cross $\gamma$ at an angle that isn’t too acute. A hyperplane $K$ is said to be $\epsilon$-almost orthogonal to $\gamma$ if $\gamma$ crosses $K$ at an angle of at least $\epsilon$.
The next lemma shows that many of the hyperplanes crossed by $\gamma$ will be almost orthogonal, for a suitable value of $\epsilon$. The proof uses elementary Euclidean trigonometry.
Lemma 3.7.
Let $d\geq 2$, and let $\gamma$ be an $\ell^{2}$ geodesic in a $d$-dimensional CAT(0) cube complex $\widetilde{X}$. Suppose that $\gamma$ crosses two hyperplanes $K_{1}$ and $K_{2}$ consecutively. Then at least one of $K_{1}$ and $K_{2}$ is $\epsilon_{d}$-almost orthogonal to $\gamma$, where $\epsilon_{d}=\tan^{-1}\left(\frac{1}{2\sqrt{d-1}}\right)$.
Proof.
If $\gamma$ crosses $K_{1}$ and $K_{2}$ simultaneously then the result is immediate, because $K_{1}$ and $K_{2}$ are orthogonal to each other and $\epsilon_{d}<\pi/4$.
Otherwise, suppose that $\gamma(0)\in K_{1}$ and $\gamma$ crosses $K_{1}$ at an angle $\theta<\epsilon_{d}$. Because the (closed) regular neighbourhood $N(K_{1})$ of $K_{1}$ is isometric to $K_{1}\times[-1/2,1/2]$, we have $\gamma(t)\in N(K_{1})$ whenever $|t|\sin\theta\leq 1/2$. In particular, if $\gamma(t_{0})$ is in the boundary of $N(K_{1})$ then $|t_{0}|=1/(2\sin\theta)$. Let $p$ be the orthogonal projection of $\gamma(t_{0})$ to $K_{1}$. The three points $\gamma(0)$, $\gamma(t_{0})$, $p$ together form a right-angled Euclidean triangle, so
$d(\gamma(0),p)=|t_{0}|\cos\theta=\frac{1}{2}\cot\theta\,.$
Because $\theta<\epsilon_{d}$, it follows that
$d(\gamma(0),p)>\frac{1}{2}\cot\epsilon_{d}=\sqrt{d-1}$
which implies that $\gamma(0)$ and $p$ are separated by a hyperplane in the $(d-1)$-dimensional cube complex $K_{1}$. Since every such hyperplane extends uniquely to a hyperplane in $\widetilde{X}$ orthogonal to $K_{1}$, it follows that $\gamma$ crosses a hyperplane in $N(K_{1})$ orthogonal to $K_{1}$.
The first such hyperplane is by assumption $K_{2}$. Let $\gamma(t_{1})\in K_{2}$ and let $q$ be the point of intersection between $K_{1}$ and $K_{2}$. Then the three points $\gamma(0)$, $\gamma(t_{1})$, $q$ together span a right-angled Euclidean triangle, whence $\gamma$ makes an angle
$\pi/2-\theta>\pi-\epsilon_{d}>\pi/4>\epsilon_{d}$
with $K_{2}$, as required. ∎
The next lemma guarantees that a sufficiently distant pair of hyperplanes will be separated by many mutually disjoint hyperplanes, all themselves almost orthogonal to a geodesic.
Lemma 3.8.
Let $H_{1}$ and $H_{2}$ be hyperplanes in a $d$-dimensional CAT(0) cube complex $\widetilde{X}$, and let $\gamma$ be an $\ell^{2}$-geodesic between them. For every $n$ there is a $C$ such that, if the length of $\gamma$ is greater than $C$, then there exists a sequence of mutually disjoint hyperplanes
$K_{1},\ldots,K_{n}$
that each separate $H_{1}$ from $H_{2}$, all of which are $\epsilon_{d}$-almost orthogonal to $\gamma$.
Proof.
Let $N$ be the Ramsey number $R(n,d+1)$. Recall that the $\ell^{1}$ distance between $H_{1}$ and $H_{2}$ is equal to the number of hyperplanes crossed by $\gamma$. Since the $\ell^{1}$ and $\ell^{2}$ metrics are quasi-isometric, there is a $C$ such that, if $\gamma$ has $\ell^{2}$ length at least $C$, then it crosses at least $2N$ hyperplanes. By Lemma 3.7, at least $N$ of these hyperplanes are $\epsilon_{d}$-almost orthogonal to $\gamma$. Since no set of $d+1$ hyperplanes of $\widetilde{X}$ pairwise intersect, it follows from the definition of $N$ that there is a pairwise disjoint set
$\{K_{1},\ldots,K_{n}\}$
of $n$ hyperplanes that intersect $\gamma$ $\epsilon_{d}$-almost orthogonally.
Finally, note that any hyperplane $K$ that crosses $H_{i}$ meets it orthogonally, and therefore if $\gamma$ also crossed $K$ there would be a compact geodesic triangle in $\widetilde{X}$ with two right angles, contradicting the CAT(0) condition. Therefore, each $K_{j}$ is disjoint from $H_{1}$ and $H_{2}$, and therefore separates them from each other. ∎
We needed to find almost orthogonal hyperplanes because of the following lemma, which will enable us to bound the diameter of a hyperplane intersecting a suitably chosen 2-cut.
Lemma 3.9.
Suppose that an $\ell^{2}$ geodesic $\gamma$ intersects a hyperplane $H$ at an angle $\theta$. Let $x$ be the point of intersection and let $y$ be a point in $H$. Then
$d(x,y)\sin\theta\leq d(y,\gamma)\,.$
In particular, if $H$ is $\epsilon$-almost orthogonal to $\gamma$, $d(x,y)\leq d(y,\gamma)/\sin\epsilon$.
Proof.
Let $z$ be the closest point on $\gamma$ to $y$, and consider the geodesic triangle with vertices $x$, $y$ and $z$. Let $\bar{\theta}$ be the angle in the Euclidean comparison triangle at the vertex corresponding to $x$, which is at least $\theta$ by [8, Proposition 1.7(4)]. Now, the Euclidean sine rule gives
$d(y,z)\geq d(x,y)\sin\bar{\theta}\geq d(x,y)\sin\theta$
as required. ∎
The proof of the claimed bound on the widths of 2-cuts uses the pigeonhole principle, applied to isomorphism types of abstract hyperplane components of a given diameter.
Remark 3.10.
Let $X$ be a compact, non-positively curved cube complex, and let $R\geq 0$. The set of isomorphism types of abstract bounding hyperplane components $Z$ over $X$ of diameter at most $R$ is finite.
This finiteness is the basis for our argument using the pigeonhole principle.
Theorem 3.11.
Let $X$ be an essential, $d$-dimensional, hyperbolic cube complex without 0-cuts or 1-cuts. Then $\pi_{1}(X)$ is cyclically indecomposable if and only if there is a bound on the widths of 2-cuts in $X$.
Proof.
The ‘only if’ direction is similar to the proof of Theorem 3.4. If $\pi_{1}(X)$ has a cyclic splitting then, by Lemma 3.3, there is an infinite cyclic subgroup $\langle\phi\rangle$ acting cocompactly on some isometrically embedded subcomplex $\widetilde{Z}\subseteq\widetilde{X}$ with $\mathrm{Wh}_{X}(\widetilde{Z})$ disconnected. As in the proof of Theorem 3.4, we can find a hyperplane $H$ of $\widetilde{Z}$ such that, for any $n$, cutting $\widetilde{Z}$ along $H$ and $\phi^{n}(H)$ gives a 2-cut between $H$ and $\phi^{n}(H)$. Since $d(H,\phi^{n}(H))\to\infty$ with $n$, it follows that there is no bound on the widths of 2-cuts, as required.
For the converse, let $Y$ be a 2-cut between hyperplanes $H_{1}$ and $H_{2}$. By Lemma 3.6, we may replace $Y$ by its intersection with $\operatorname{Hull}(H_{1}\cup H_{2})$, and so assume that $Y$ is contained in $\operatorname{Hull}(H_{1}\cup H_{2})$. Let $\gamma$ be an $\ell^{2}$ geodesic between $H_{1}$ and $H_{2}$. It is easy to see that $H_{1}\cup\gamma\cup H_{2}$ is $2\delta$-quasiconvex in the $\ell^{2}$ metric and therefore, by Theorem 1.7, $\operatorname{Hull}(H_{1}\cup H_{2})$ is contained in the $R$-neighbourhood of $H_{1}\cup\gamma\cup H_{2}$, for some uniform $R$.
Let $M$ be the number of isomorphism types of oriented abstract hyperplane components over $X$ of diameter at most $2R\sqrt{4d-3}$ (which is finite by Remark 3.10). Let $M^{\prime}$ be such that any geodesic crossing $M^{\prime}$ hyperplanes has an $\ell^{2}$ length of at least $(1+\sqrt{4d-3})R$, and let $N=M+2M^{\prime}$. By Lemma 3.8, if the width of $Y$ is greater than a constant $C$, then we can find a sequence of $M+1$ hyperplanes
$K_{0},K_{1},\ldots,K_{M}\,,$
all separating $H_{1}$ from $H_{2}$, all $\epsilon_{d}$-almost orthogonal to $\gamma$, and such that each $K_{i}$ intersects $\gamma$ at distance greater than $(1+\sqrt{4d-3})R$ from both $H_{1}$ and $H_{2}$.
From each $K_{i}$ we define an oriented abstract hyperplane component $Z_{i}$ as follows. First, set $Z_{i}=K_{i}\cap Y$; we next need to partition $\pi_{0}(\mathrm{Wh}_{K_{i}}(Z_{i}))$. Cutting $Y$ along $K_{i}$ decomposes it as
$Y=Y_{1}\cup_{K_{i}}Y_{2}\,,$
and Proposition 2.21 then tells us that
$\mathrm{Wh}_{X}(Y)=\mathrm{Wh}_{X}(Y_{1})\#_{K_{i}}\mathrm{Wh}_{X}(Y_{2})\,.$
In particular, $\mathrm{Wh}_{X}(Y)$ contains a copy of $\mathrm{Wh}_{K_{i}}(Z_{i})$, which is canonically isomorphic to the links of $K_{i}$ in both $\mathrm{Wh}_{X}(Y_{i})$ by Lemma 2.13. The partition on $\pi_{0}(\mathrm{Wh}_{K_{i}}(Z_{i}))$ is now induced by connectivity in $\mathrm{Wh}_{X}(Y)\smallsetminus\{H_{1},H_{2}\}$. Furthermore, each $Z_{i}$ is naturally oriented from $H_{1}$ to $H_{2}$.
We next prove a bound on the diameter of any $Z_{i}$. Indeed, let $x$ be the point at which $K_{i}$ intersects $\gamma$. Since $Y$ is contained in the closed $R$-neighbourhood of $H_{1}\cup\gamma\cup H_{2}$, $Z_{i}$ is the union of two closed subsets:
$Z_{i}=P\cup Q$
where $P=Z_{i}\cap(N_{R}(H_{1})\cup N_{R}(H_{2}))$ and $Q=Z_{i}\cap N_{R}(\gamma)$.
By Lemma 3.9, the distance from any $y\in Q$ to $x$ is bounded above by
$\frac{R}{\sin\epsilon_{d}}=\frac{R}{\sin\cot^{-1}(2\sqrt{d-1})}=R\sqrt{4d-3}\,.$
We next argue that $P$ is empty. Indeed, if not then there is $y\in P\cap Q$, since $Z_{i}$ is connected and $P$ and $Q$ are both closed. But such a $y$ is within distance $R$ of $H_{1}\cup H_{2}$ and within $R\sqrt{4d-3}$ of $x$, so $K_{i}$ intersects $\gamma$ at distance less than $R(1+\sqrt{4d-3})$ from either $H_{1}$ or $H_{2}$, contradicting the construction of the $K_{i}$ above. Thus, $P$ is indeed empty, so $Z_{i}=Q$ and $\operatorname{diam}(Z_{i})=\operatorname{diam}(Q)\leq 2R\sqrt{4d-3}$.
Therefore, by the pigeonhole principle and the choice of $M$, there exist distinct $i<j$ such that $Z_{i}$ and $Z_{j}$ are isomorphic oriented abstract hyperplane components. Cutting $Y$ along $K_{i}$ and $K_{j}$ decomposes it as
$Y=Y_{1}\cup_{K_{i}}Y^{\prime}\cup_{K_{j}}Y_{2}\,.$
Applying Lemma 2.24 twice, $Y^{\prime}$ is a 2-cut between $K_{i}$ and $K_{j}$, while $Z_{i}$ and $Z_{j}$ are the corresponding bounding hyperplane components. Therefore, $Y^{\prime}$ is in fact a periodic 2-cut between $K_{i}$ and $K_{j}$, as required. ∎
3.3 From $2$-cuts to $k$-cuts
In the previous section, we proved a uniform bound on the widths of 2-cuts. In this section, we will upgrade this to a uniform bound on the widths of $k$-cuts, where the bound is independent of $k$. The key result is Lemma 3.16, which encapsulates the fact that a finite collection of hyperplanes coarsely has the structure of a quasi-tree. I am grateful to Mark Hagen for pointing out that this lemma should be true, and that it would simplify the proof my argument.
The proof of Lemma 3.16 is inspired by the projection-complex machinery of Bestvina–Bromberg–Fujiwara [5]. Consider a collection $\mathbb{V}=\{V_{i}\}$ of convex subcomplexes of a hyperbolic, CAT(0) cube complex $\widetilde{X}$. (In [5] this collection would typically be infinite. For us it will be finite, but of unbounded cardinality.) For each $U\in\mathbb{V}$, we consider the closest-point projection $\pi_{U}:\widetilde{X}\to U$. For each $U$, Bestvina–Bromberg–Fujiwara use this to define a distance function on pairs of elements of $\mathbb{V}\smallsetminus\{U\}$ via
$d^{\pi}_{U}(V,W):=\operatorname{diam}(\pi_{U}(V)\cup\pi_{U}(W))\,.$
They go on to state projection complex axioms for the distance functions $d^{\pi}_{\bullet}$, with respect to a fixed constant $\theta>0$; see [6, §1.3] for a concise statement of the axioms. The next proposition asserts that, as long as the elements of $\mathbb{V}$ are pairwise sufficiently distant from each other, the projection complex axioms are satisfied.
Proposition 3.12.
Let $\widetilde{X}$ be a $\delta$-hyperbolic CAT(0) cube complex and let $\mathbb{V}$ be a finite set of convex subcomplexes of $\widetilde{X}$. There is a constant $C$, depending only on $\widetilde{X}$, such that, as long as the elements of $\mathbb{V}$ are pairwise at distance at least $C$, there is $\theta>0$ such that the following hold for all distinct $U,V,W,Y\in\mathbb{V}$:
  1. (i)
    $d^{\pi}_{U}(V,W)=d^{\pi}_{U}(W,V)$;
  2. (ii)
    $d^{\prime\pi}_{U}(V,W)+d^{\pi}_{U}(W,Y)\geq d^{\pi}_{U}(V,Y)$;
  3. (iii)
    $d^{\pi}_{U}(V,V)\leq\theta$;
  4. (iv)
    if $d^{\pi}_{U}(V,W)>\theta$ then $d_{V}(U,W)\leq\theta$;
  5. (v)
    #$\#\{Z\in\mathbb{V}\mid d^{\pi}_{Z}(V,W)>\theta\}<\infty$
Proof.
Properties (i) and (ii) are obvious, while property (v) is immediate by hypothesis since $\mathbb{V}$ is assumed to be finite. Property (iii) is known as the contracting property for convex subspaces of $\delta$-hyperbolic spaces, and a standard exercise shows that it holds for any $\theta\geq 8\delta$ as long as $C\geq 2\delta$. Finally, property (iv) is known as the Behrstock inequality. A sketch is given in [44, p. 133]. ∎
In order to construct a quasitree, it is necessary to adjust the distance functions $d^{\pi}_{\bullet}$ to strengthen the Behrstock inequality. For this, we follow Bestvina–Bromberg–Fujiwara–Sisto [6]. From $d^{\pi}_{\bullet}$, they construct a new family of distance functions $d_{\bullet}$ with improved properties.
Proposition 3.13.
For $d^{\pi}_{\bullet}$ and $\theta$ as above, there are functions $d_{\bullet}$ such that
$|d^{\pi}_{\bullet}-d_{\bullet}|\leq 2\theta\,,$
and that satisfy the following axioms for all distinct $U,V,W,Y\in\mathbb{V}$:
  1. (i)
    $d_{U}(V,W)=d_{U}(W,V)$;
  2. (ii)
    $d_{U}(V,W)+d_{U}(W,Y)\geq d_{U}(V,Y)$;
  3. (iii)
    $d_{U}(V,V)\leq 11\theta$;
  4. (iv)
    if $d_{U}(V,W)>11\theta$ then $d_{W}(V,Y)=d_{W}(U,Y)$ unless $Y=W$;
  5. (v)
    #$\#\{Z\in\mathbb{V}\mid d_{Z}(V,W)>\theta\}<\infty$.
Proof.
This is [6, Proposition 4.14]. ∎
Remark 3.14.
Proposition 3.13(iv) should be thought of as a strengthened version of the Behrstock inequality. Indeed, taking $Y=V$ in Proposition 3.13(iv) shows that, if $d_{U}(V,W)>11\theta$, then $d_{W}(U,V)=d_{W}(V,V)\leq 11\theta$ by Proposition 3.13(iii).
Bestvina–Bromberg–Fujiwara–Sisto use the strengthened projection complex axioms to construct a quasitree. Instead, we will use them to find $U\in\mathbb{V}$ onto which the diameter of the projection of all the other elements of $\mathbb{V}$ is bounded.
Lemma 3.15.
Let $\widetilde{X}$ be a $\delta$-hyperbolic CAT(0) cube complex and let $\mathbb{V}$ be a finite set of unbounded convex subcomplexes of $\widetilde{X}$. There is a constant $C$, depending only on $\widetilde{X}$, such that, as long as the elements of $\mathbb{V}$ are pairwise at distance at least $C$, there is $U\in\mathbb{Y}$ such that
$\operatorname{diam}\left(\bigcup_{V\in\mathbb{V}\smallsetminus U}\pi_{U}(V)\right)\leq 13\theta\,.$
Proof.
Take $C$ as in Proposition 3.12. Since $d_{\bullet}$ is within $2\theta$ of $d^{\pi}_{\bullet}$, it suffices to prove that there is $U$ such that $d_{U}(V,W)\leq 11\theta$ for all $V,W\neq U$. We prove this by induction on the cardinality of $\mathbb{V}$.
The base cases when #$\#\mathbb{V}\leq 2$ are trivial. For the inductive step, fix $W_{0}\in\mathbb{V}$ and let $\mathbb{V}^{\prime}:=\mathbb{V}\smallsetminus\{W_{0}\}$. By induction, there is $U^{\prime}\in\mathbb{V}^{\prime}$ such that $d_{U^{\prime}}(V^{\prime},W^{\prime})\leq 11\theta$ for all $V^{\prime},W^{\prime}\in\mathbb{V}^{\prime}\smallsetminus\{U^{\prime}\}$.
If $d_{U^{\prime}}(V^{\prime},W_{0})\leq 11\theta$ for all $V^{\prime}\in\mathbb{V}^{\prime}$ then we are done, taking $U=U^{\prime}$. Suppose, therefore, that $d_{U^{\prime}}(V^{\prime},W_{0})>11\theta$ for some $V^{\prime}\in\mathbb{V}^{\prime}\smallsetminus\{U^{\prime}\}$. In this case, we set $U:=W_{0}$, and we need to show that $d_{U}(V,W)\leq 11\theta$ for any $V,W\in\mathbb{V}^{\prime}$.
Proposition 3.13(iv) implies that
$d_{U^{\prime}}(V,U)=d_{U^{\prime}}(V,W_{0})=d_{U^{\prime}}(V^{\prime},W_{0})>11\theta$
for any $V\neq U^{\prime}$, while $d_{U^{\prime}}(U^{\prime},U)>11\theta$ since $U^{\prime}$ is unbounded. Applying Proposition 3.13(iv) again, it follows that $d_{U}(V,W)=d_{U}(V,U^{\prime})$ for any $W\neq U$. But $d_{U}(V,U^{\prime})\leq 11\theta$ by Remark 3.14, so $d_{U}(V,W)\leq 11\theta$ for any $V,W\neq U$, as required. ∎
The geometric lemma that we need can now be easily deduced.
Lemma 3.16.
Let $\widetilde{X}$ be a $d$-dimensional, $\delta$-hyperbolic CAT(0) cube complex. There is a constant $D$ with the following property. Suppose that $\mathbb{V}$ is any finite, non-empty set of unbounded convex subcomplexes of $\widetilde{X}$. If $d(U,V)>D$ for every distinct $U,V\in\mathbb{V}$, then there is a hyperplane $H$ and $U\in\mathbb{V}$ such that $H$ separates $U$ from every $V\in\mathbb{V}\smallsetminus\{U\}$.
Proof.
Let $C$ and $\theta$ be as in Lemma 3.15. By Theorem 1.7, there is $r$ such that, for any subset $Z\subseteq\widetilde{X}$ of diameter at most $13\theta$, the diameter of $\operatorname{Hull}(Z)$ is at most $r$.
For each $U\in\mathbb{V}$, let $\widehat{U}$ be the convex hull of the $(\sqrt{d}+r)$-neighbourhood of $U$, which is contained in the $R$-neighbourhood of $U$, for some uniform $R$, by Theorem 1.7. Let $\widehat{\mathbb{V}}:=\{\widehat{U}\mid U\in\mathbb{V}\}$ and let $D=C+2R$.
As long as the elements of $\mathbb{V}$ are at distance at least $D$ from each other, it follows that the elements of $\widehat{\mathbb{V}}$ are at distance at least $C$ from each other. Hence, by Lemma 3.15, there is $\widehat{U}$ such that
$\operatorname{diam}\left(\bigcup_{\widehat{V}\in\widehat{\mathbb{V}}\smallsetminus\widehat{U}}\pi_{\widehat{U}}(\widehat{V})\right)\leq 13\theta\,,$
whence
$\operatorname{diam}\operatorname{Hull}\left(\bigcup_{\widehat{V}\in\widehat{\mathbb{V}}\smallsetminus\widehat{U}}\pi_{\widehat{U}}(\widehat{V})\right)\leq r$
by the definition of $r$.
This hull contains a point on the frontier of $\widehat{U}$, so is at distance at least $\sqrt{d}$ from $U$ itself. Therefore, any shortest path from $U$ to the hull crosses some hyperplane $H$. Thus, $H$ separates $U$ from $\pi_{\widehat{U}}(V)$, for every $V\neq U$. It follows that $H$ separates $U$ from $V$, for all $V\neq U$, as required. ∎
We can now upgrade our bound on the widths of 2-cuts to a bound on the widths of $k$-cuts.
Theorem 3.17.
Let $X$ be an essential, $d$-dimensional, hyperbolic cube complex without 0-cuts or 1-cuts. Then $\pi_{1}(X)$ is cyclically indecomposable if and only if there is a uniform bound, independent of $k$, on the widths of all $k$-cuts with $k\geq 2$.
Proof.
If there is a bound on the widths of $k$-cuts for all $k\geq 2$ then, in particular, there is a bound on the widths of all $2$-cuts. Therefore, $X$ is cyclically essential and $\pi_{1}(X)$ is cyclically indecomposable by Theorem 3.11.
For the converse implication, let $C$ be the bound from Theorem 3.11. By Theorem 1.7, there is $R$ such that the hull of the $C$-neighbourhood of any convex subset is contained in the $R$ neighbourhood of that subset. Let $D$ be the constant from Lemma 3.16.
Let $Y$ be a $k$-cut between a collection of $k\geq 2$ hyperplanes $\{H_{1},\ldots H_{k}\}$, and suppose that the width of $Y$ is greater than $2R+D$. Suppose, furthermore, that $k$ is minimal such that there exists a $k$-cut of width greater than $2R+D$.
For each $i$, let $U_{i}$ be the convex hull of the $C$-neighbourhood of $H_{i}$.Then the $U_{i}$ are pairwise at distance at least $D$.
By Lemma 3.16, there is a hyperplane $K$ that, without loss of generality, separates $U_{1}$ from the remaining $H_{i}$. In particular, $K$ is at distance at least $C$ from $H_{1}$ and separates all the other $H_{i}$ from $H_{1}$. Thus, cutting $Y$ along $K$ gives
$Y=Y_{1}\cup_{K}Y_{2}$
where $H_{1}$ bounds $Y_{1}$ and the remaining $H_{i}$ all bound $Y_{2}$.
By the minimality of $k$, $\{H_{2},\ldots,H_{k}\}$ is not a cut set in $\mathrm{Wh}_{X}(Y_{2})$. Therefore, Lemma 2.24 implies that $Y_{1}$ is a 2-cut between $H_{1}$ and $K$, which is of width at least $C$ by construction. By the choice of $C$, it follows that $\pi_{1}(X)$ is cyclically decomposable, as required. ∎
4 One-ended subgroups
In this section, we prove both our main theorem about cubulated hyperbolic groups and its corollary about one-relator groups.
4.1 Cubulated hyperbolic groups
With the uniform bound on the widths of cuts in hand, it is easy to prove the main theorem in the cyclically indecomposable case.
Theorem 4.1.
Let $X$ be a hyperbolic cube complex. If $\pi_{1}(X)$ is one-ended and cyclically indecomposable then it has a one-ended quasiconvex subgroup of infinite index.
Proof.
By Corollary 2.32, we may without loss of generality assume that $X$ is essential, without 0- or 1-cuts. Let $C$ be the uniform bound on widths of $k$-cuts given by Theorem 3.17.
Choose any hyperplane $H$ of $X$, and fix an elevation $\widetilde{H}$ to the universal cover $\widetilde{X}$. Because $\pi_{1}H$ is separable in $G=\pi_{1}X$ by the work of Agol [1] and Haglund–Wise [26], there is a finite-index subgroup $G_{0}$ of $G$ such that
$d(\widetilde{H},g\widetilde{H})>C$
for any $g\in G_{0}\smallsetminus\operatorname{Stab}(\widetilde{H})$.
Let $X_{0}$ denote the quotient $G_{0}\backslash\widetilde{X}$ and let $H_{0}$ be the image of $\widetilde{H}$, which is a hyperplane of $X_{0}$. Let $X^{\prime}$ be (a component of) the result of cutting $X_{0}$ along $H_{0}$. Since $X^{\prime}$ is finite and maps locally isometrically to $X$, $\pi_{1}(X^{\prime})$ is a quasiconvex subgroup of $G$. Furthermore, since $\pi_{1}(X^{\prime})$ is a vertex-group of a non-trivial graph-of groups decomposition for $G_{0}$, it has infinite index in $G_{0}$, and hence in $G$. It remains to show that $\pi_{1}(X^{\prime})$ is one-ended.
With a view to a contradiction, suppose that $\pi_{1}(X^{\prime})$ is freely decomposable. By Proposition 2.18, there is a 0-cut $Y\to X^{\prime}$. But $\mathrm{Wh}_{X^{\prime}}(Y)$ is obtained from $\mathrm{Wh}_{X}(Y)$ by deleting the $G_{0}$-translates of $\widetilde{H}$, so some finite set of translates
$g_{1}\widetilde{H},\ldots,g_{k}\widetilde{H}$
(with $g_{i}\in G_{0}$) define a cut set in $\mathrm{Wh}_{X}(Y)$. Since $\mathrm{Wh}_{X}(Y)$ has no 0-cut or 1-cuts, $k\geq 2$. Therefore $Y$ is a $k$-cut in $\widetilde{X}$ of width greater than $C$, contradicting the choice of $C$. ∎
By passing to the Grushko and Bowditch JSJ decompositions, the main theorem follows.
Proof of Theorem A.
Since $\pi_{1}(X)$ is a finitely generated, torsion-free group, it is freely indecomposable if and only if it is one-ended, by Stallings’ theorem [45]. If $\pi_{1}(X)$ is freely decomposable, Grushko’s theorem implies that either $\pi_{1}(X)$ is free or it has a one-ended free factor, which is necessarily a quasiconvex subgroup of infinite index. Therefore, we may assume that $\pi_{1}(X)$ is itself one-ended.
Consider the Bowditch JSJ decomposition of $\pi_{1}(X)$ [7]. If the decomposition has only one vertex and no edges, then that vertex is either of surface or of rigid type. In the former case, $\pi_{1}(X)$ is a surface group. In the latter case, $\pi_{1}(X)$ is cyclically indecomposable, so it has a one-ended, quasiconvex subgroup of infinite index by Theorem 4.1.
Now consider the case where the JSJ decomposition is non-trivial. Every vertex group $V$ is quasiconvex [7, Proposition 1.2], and in particular finitely generated and hyperbolic. If some vertex group $V$ is not itself free, then by Grushko’s theorem it has a one-ended free factor $V^{\prime}$, which provides a one-ended, quasiconvex subgroup, necessarily of infinite index in $\pi_{1}(X)$. On the other hand, if every vertex group is free, then $\pi_{1}(X)$ is either a surface group or contains a one-ended, quasiconvex subgroup by [52, Theorem 3]. ∎
4.2 One-relator groups
Combining the result for cubulated hyperbolic groups with recent joint work of the author with Louder [30, 31], and also work of Linton [29], we can also answer Question 0.1 for one-relator groups. The key is Puder’s notion of primitivity rank [37].
Definition 4.2 (Primitivity rank).
Let $w$ be a non-trivial element of a free group $F$. The primitivity rank $\pi(w)$ is the minimal rank of a subgroup $w\in H\subseteq F$ such that $w$ is not a primitive element of $H$.
By convention, $\min\varnothing=\infty$, so $\pi(w)=\infty$ if and only if $w$ is primitive in $F$ (and hence in every subgroup containing it). As a result of this convention, $\pi(w)\geq 1$ for all $w$. (It can be convenient to extend the definition of $\pi$ to the trivial element $1\in F$ and set $\pi(1)=0$.)
[…]
Theorem 4.3.
[…]
  1. (i)
    $\pi(w)=\infty$ if and only if $G$ is free;
  2. (ii)
    $\pi(w)=1$ if and only if $G$ has torsion;
  3. (iii)
    if $\pi(w)=2$ then $G$ has a two-generator, one-relator subgroup $P$ that is not free;
  4. (iv)
    if $\pi(w)>2$ then $G$ is hyperbolic and cubulated.
Proof.
Item (i) is a result of the classical theorem of Whitehead that, for non-trivial $w$, $G$ is free if and only if $w$ is primitive [32, Proposition 5.10]. It is easy to see that $\pi(w)=1$ if and only if $w$ is a proper power in $F$; a classical theorem of Karrass–Magnus–Solitar asserts that this is true exactly when $G$ has torsion [32, Proposition 5.17]. Item (iii) is a theorem of Louder and the author [30, Theorem 1.5], while item (iv) is a theorem of Linton [29, Theorem 8.2] (using deep work of Wise [56], as well as further work of Louder and the author [31]). ∎
Remark 4.4.
The one-relator group $G$ is also hyperbolic and cubulated when $\pi(w)=1$, by work of Newman [35] and Wise [56]. However, we will not need either of these facts here.
The final piece of the jigsaw is a recent result of Gardam–Kielak–Logan, who handled the two-generator, one-relator case of Question 0.1 [20, Theorem 1.5].
Theorem 4.5 (Gardam–Kielak–Logan).
Let $G$ be a two-generator, one-relator group. If every subgroup of infinite index in $G$ is free, then $G$ is either free or the fundamental group of a closed surface.
Combining these results with Theorem A, we obtain the desired result for all one-relator groups.
Proof of Theorem D.
[…]
In this case, $G$ has a two-generator, one-relator subgroup $P$ that is not free by Theorem 4.3(iii). The index of $P$ in $G$ is finite by hypothesis, and every subgroup of infinite index in $P$ is free. By Theorem 4.5, $P$ is either $\mathbb{Z}^{2}$ or $\mathbb{Z}\rtimes\mathbb{Z}$, so the same is true of $G$ by Bieberbach’s theorem. ∎
5 Questions
This final section contains some open questions that are suggested by the results of this paper.
Strebel’s theorem asserts that every subgroup of infinite index in an $n$-dimensional Poincaré duality group has cohomological dimension at most $n-1$ [48]. Since groups of cohomological dimension one are precisely free groups [49], Theorems A and D can be thought of as converses to Strebel’s theorem for dimension $n=2$, in the contexts of cubulated hyperbolic and one-relator groups. This is the whole story for one-relator groups, which are always virtually of cohomological dimension 2. However, for cubulated hyperbolic groups, it makes sense to ask whether the converses of the higher-dimensional instances of Strebel’s theorem are also true.
Question 5.1.
Let $X$ be a hyperbolic cube complex with $\operatorname{cd}(\pi_{1}(X))=n>2$. If the cohomological dimension of every subgroup of infinite index in $\pi_{1}(X)$ is less than $n$, does it follow that $\pi_{1}(X)$ must be a $PD_{n}$ group?
As with Theorem A, it is also conceivable that the geometric hypotheses of Question 5.1 can be relaxed.
The proof of Theorem A made essential use of the separability of hyperplane stabilisers, which is closely related to the fact that $X$ is virtually special. Removing this condition would necessitate a different approach.
Question 5.2.
Let $X$ be a compact, non-positively curved cube complex with every subgroup of infinite index free. Must $\pi_{1}(X)$ be either a free group or a surface group?
An answer to Question 5.2 in the case of lattices in products of trees would be interesting. The well known flat-closing conjecture would imply a positive answer.
In a similar vein, removing the cubulation hypothesis from Theorem A would be a major breakthrough.
Question 5.3.
Let $G$ be an infinite hyperbolic group. If every subgroup of infinite index in $G$ is free, must $G$ be either a free group or a surface group?
Of course, if every one-ended hyperbolic group had a surface subgroup, then Question 5.3 would have an affirmative answer.
The next question addresses the slight mismatch between the conclusions of our main theorems.
Question 5.4.
Let $G$ be a two-generator, one-relator group such that every finitely generated subgroup of infinite index is free. Must $G$ be isomorphic to either $\mathbb{Z}$ or a solvable Baumslag–Solitar group?
(Note that the two two-generator surface groups, $\mathbb{Z}^{2}$ and $\mathbb{Z}\rtimes\mathbb{Z}$, are $BS(1,1)$ and $BS(1,-1)$ respectively.)
As explained in §2.5, any non-positively curved square complex can be ‘unfolded’ until Theorem E applies, providing an effective geometric computation of the Grushko decomposition. I would like to extend this method to higher-dimensional cube complexes. [^5] Touikan’s algorithm provides a general computation of the Grushko decomposition for any torsion-free group with solvable word problem [50]. However, a cubical method would be preferable.
The first step is to recognise whether or not the cube complex $X$ is essential. Sam Shepherd has explained to me that being essential can be recognised algorithmically: the key observation is that an inessential CAT(0) cube complex contains a half-space that is minimal with respect to inclusion, and the bounding hyperplane of a minimal half-space can be recognised [43].
Given essential $X$, the next step is to determine whether or not $\pi_{1}(X)$ splits freely. The following question asks whether or not the sufficient condition of Theorem E can always be applied.
Question 5.5.
Let $G$ be a one-ended cubulable group. Does $G=\pi_{1}(X)$, where $X$ is essential and no vertex of $X$ has a link with a separating simplex?
Shepherd’s theorem (Theorem 2.30) implies that there is an $X$ without separating vertices in the links, but does not provide information about higher-dimensional simplices. Regardless of the answer to Question 5.5, it is also important to find an algorithmic condition to recognise whether or not $X$ admits 0- or 1-cuts.
Assuming an understanding of free splittings, one may pose similar questions about splittings over $\mathbb{Z}$. It is not clear if there is an analogue of Theorem E for cyclic splittings.
Question 5.6.
Let $X$ be a compact, non-positively curved cube complex with $\pi_{1}(X)$ one-ended. Is there a condition on the links of vertices of $X$ that guarantees that $\pi_{1}(X)$ does not split over $\mathbb{Z}$?
For instance, it seems probable that the techniques of this article can be used to prove that, if $X$ is a hyperbolic square complex and the links of vertices of $X$ are connected without cut vertices or cut pairs, then $\pi_{1}(X)$ does not split over $\mathbb{Z}$. But it would be better still to provide a local criterion for higher-dimensional cube complexes, or outside the hyperbolic setting.
As a step towards understanding cyclic splittings outside the hyperbolic setting, the next question asks whether the hyperbolicity hypothesis can be removed from Theorem 3.4.
Question 5.7.
Let $X$ be a compact, essential, non-positively curved cube complex without 0- or 1-cuts. If $\pi_{1}(X)$ splits over $\mathbb{Z}$, must $X$ admit a periodic 2-cut?
The converse statement, that $\pi_{1}(X)$ splits over $\mathbb{Z}$ if $X$ admits a periodic 2-cut, follows from the algebraic annulus theorem of Dunwoody–Swenson [17].